### v1

# Minimal Construction of Graphs with Maximum Robustness

Abstract
The notions of network $r$-robustness and $(r,s)$-robustness have been earlier introduced in the literature to achieve resilient control in the presence of misbehaving agents. However, while higher robustness levels provide networks with higher tolerances against the misbehaving agents, they also require dense communication structures, which are not always desirable for systems with limited capabilities and energy capacities. Therefore, this paper studies the fundamental structures behind $r$-robustness and $(r,s)$- robustness properties in two different ways. (a) We first explore and establish the tight necessary conditions on the number of edges for undirected graphs with any nodes must satisfy to achieve maximum $r$- and $(r,s)$-robustness. (b) We then use these conditions to construct two classes of undirected graphs, referred as to $\gamma$- and $(\gamma,\gamma)$-Minimal Edge Robust Graphs (MERGs), that provably achieve maximum robustness with minimal numbers of edges. We finally validate our work through some sets of simulations.
{IEEEkeywords}
Network control systems, distributed consensus and control, resilient consensus
1 Introduction
[…]
Distributed multi-agent systems have many practical uses, ranging from information gathering [1], target tracking [2], to collaborative decision making [3]. A fundamental multi-agent systems’ objective is consensus, under which multiple agents reach agreement on common state values. However, it is known that performance of consensus deteriorates significantly when one or more misbehaving agents share wrong, or even adversarial, information with the system. Resilient consensus in the presence of misbehaving agents has been studied extensively [4, 5, 6, 7, 8, 9, 10, 11]. In [4, 6, 8], an algorithm called Weighted Mean-Subsequence-Reduced (W-MSR) was introduced to allow the non-misbehaving agents (often called normal) to reach consensus despite the presence of misbehaving agents. Unlike approaches such as those in [12, 13, 14], which rely on trust assessments - often at the cost of increased computational complexity - W-MSR instead achieves resilient consensus through a rule-based filtering strategy that discards extreme values. When applied to graphs satisfying the topological properties of $r$-robustness or $(r,s)$-robustness, W-MSR guarantees that normal agents can converge to a value within the convex hull of their initial conditions, even in the presence of a bounded number of adversaries.
In general, a higher robustness in a communication graph allows normal agents to tolerate a larger number of misbehaving agents while achieving consensus through the W-MSR algorithm [4]. Consequently, maximizing the robustness of a graph is highly desirable. However, achieving higher robustness by definition requires a higher number of edges, often referred to as edge counts. This results in dense and complex graph structures, which can pose challenges for certain systems due to practical constraints such as limited communication range, bandwidth, and energy. This clear tradeoff between robustness levels and edge counts motivates our interest in finding the fundamental structures behind these robust graphs. We first discover the smallest edge counts possible to achieve maximum levels of robustness. Leveraging this, in this paper, we then introduce and construct two classes of graphs that achieve the maximum robustness with a minimal set of edges for any given number of nodes.
Several studies have attempted to indirectly address the tradeoff between robustness and edge counts by relaxing the connectivity of graphs. Some have explored time-varying network topologies to reduce connectivity requirements [15, 16, 17]. In [18], the concept of trusted agents, who are always deemed fault-free, is introduced to maintain robustness with fewer edges. Furthermore, robustness of multi-hop communication graphs has been investigated in [19]. While these works aim to relax the connectivity requirements for $r$- and $(r,s)$-robustness to some extent, their results still remain constrained by the structures fundamental to the robustness definitions. A deeper understanding of the fundamental or minimal graph structures that preserve $r$- and $(r,s)$-robustness would further provide insights into these methods for effectively reducing these connectivity constraints.
Despite its merits, identifying the fundamental structures for given levels of robustness is particularly challenging due to its combinatorial nature. In fact, determining $r$ robustness is a coNP-complete problem [8]. Although algorithms for computing exact robustness exist [20, 21], they scale poorly with the number of nodes. This computational bottleneck makes brute-force search approaches impractical. Although there have been works to speed up computations with sampling-based [22], heuristics [23], or machine learning-based approaches [24], they either provide approximations without guarantees or require specific assumptions about graph topologies. In addition, other works under-bound a given graph’s robustness with its algebraic connectivity $\lambda_{2}$ (the second smallest eigenvalue of a graph’s Laplacian matrix) [25, 26] or with its minimum degree [27, 28], which are generally simpler and faster to compute. However, these bounds can be too loose for many graph structures, limiting its practicality in general applications.
To circumvent the combinatorial issue, many researches have focused on the construction and analysis of specific graph classes with predetermined $r$- and $(r,s)$-robustness properties. A complete graph with $n$ nodes is $\lceil\frac{n}{2}\rceil$-robust, indicating that $\lceil\frac{n}{2}\rceil$ is the maximum level of $r$-robustness any graph of $n$ nodes can achieve [4]. Authors of [29] presented a class of circulant digraphs whose robustness levels are directly related to their connectivity parameter $k$. A preferential-attachment method is investigated in [6, 4, 8] for scalable robust graph constructions. The authors in [30] study the construction of a class of undirected $r$-robust graphs with minimum number of nodes to achieve certain robustness. Furthermore, robustness levels of various graph structures, such as geometric and random graphs [31, 32, 8, 33, 34] have been explored.
Still, understanding the fundamental structural properties of general $r$- and $(r,s)$-robust graphs remains an open problem. While the authors in [35] have established the lower bounds on edge counts and construction mechanisms for undirected $(2,2)$-robust graphs with any number of nodes, these results do not generalize to other robustness levels. In [6, 4], the minimum degree requirements for $r$- and $(r,s)$-robust graphs are presented. Nevertheless, these bounds are local in the sense that they only apply to individual nodes. The concept of minimal edge sets for $(r,s)$-robust graphs has been introduced in [36], though specific scalable construction mechanisms are not discussed.
1.0.1 Contributions
In this paper, we explicitly explore the necessary structural properties of graphs with maximum $r$- and $(r,s)$-robustness, building upon our previous work [37], in which we have constructed graphs that achieve maximum $r$-robustness with a minimal set of edges. In this paper, we aim to paint the complete picture of the tradeoff between the robustness levels and edge counts by also considering $(r,s)$-robustness and presenting related analyses. In addition, we further generalize the class of graphs presented in [37] [^1] In [37], we referred to graphs that achieve maximum robustness with minimal sets of edges as $(2r-1,r)$- and $(2r,r)$-robust graphs. To avoid confusion with the standard definition of $(r,s)$-robustness, we adopt a new naming convention in this paper, detailed later in this section.. Finally, we empirically support our results through simulations by showing that removing even a single edge immediately reduces the networks’ robustness and thus the maximum number of the misbehaving agents that can be tolerated in achieving consensus through the W-MSR algorithm. In summary, we make the following contributions in this work:
  1. 1.
    We establish the necessary conditions on the edge counts, in the forms of tight lower bounds, that undirected graphs with any number of nodes must satisfy to achieve the maximum $r$- or $(r,s)$-robustness.
  2. 2.
    Based on these necessary conditions, we construct two classes of undirected graphs with a varying number of nodes, referred to as $\gamma$- and $(\gamma,\gamma)$-Minimal Edge Robust Graphs (MERGs), in the paper, that are proven to achieve maximum $r$- and $(r,s)$-robustness with minimal sets of edges, respectively. Then, we further strengthen the result by proving that the construction mechanisms for the maximum $(r,s)$-robustness are in fact necessary and sufficient.
  3. 3.
    Through simulation, we empirically demonstrate that $\gamma$- and $(\gamma,\gamma)$-MERGs maintain maximum robustness in a minimalistic manner, with the removal of any single edge resulting in a decrease in robustness.
The remainder of this paper is organized as follows. Section 2 presents preliminaries and the problem statement. Section 3 and 4 focus on necessary structures for $r$-robustness: Section 3 establishes necessary conditions on edge counts for graphs to achieve maximum $r$-robustness, whereas Section 4 provides systematic constructions of graphs that use the fewest number of edges to achieve maximum $r$-robustness. Similarly, Section 5 and 6 examine the necessary structures for maximum $(r,s)$-robustness, where Section 5 determines the necessary conditions on edge counts for graphs to achieve the maximum $(r,s)$-robustness, and Section 6 discusses how to construct graphs that achieve maximum $(r,s)$-robustness with the smallest edge counts. Section 7 shows the simulation results, and in Section 8 conclusions and discussions of future work directions are given.
2 Background and Problem Statement
2.1 Notation
We denote a simple, undirected time-invariant graph as $\mathcal{G}=(\mathcal{V},\mathcal{E})$ where $\mathcal{V}$ and $\mathcal{E}$ are the vertex set and the undirected edge set of graph respectively. We denote the number of nodes $n$. An undirected edge $(i,j)\in\mathcal{E}$ indicates that information is exchanged between nodes $i$ and $j$. The neighbor set of agent $i\in\mathcal{V}$ is denoted as $\mathcal{N}_{i}=\{j\in\mathcal{V}\;|\;(i,j)\in\mathcal{E}\}$. The state of agent $i$ at time $t$ is denoted as $x_{i}[t]$. We denote the cardinality of a set $\mathcal{S}$ as $|\mathcal{S}|$. We denote the set of non-negative and positive integers as $\mathbb{Z}_{\geq 0}$ and $\mathbb{Z}_{>0}$.
2.2 Weighted Mean Subsequence Reduced Algoritheorem
In this section, we review the relevant fundamental concepts of $r$- and $(r,s)$-robustness of a graph, as well as the W-MSR algorithm, which are defined in [4].
Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be a graph. Then, an agent $i\in\mathcal{V}$ shares its state $x_{i}[t]$ at time $t$ with all neighbors $j\in\mathcal{N}_{i}$, and each agent $i$ updates its state according to the nominal update rule:
$x_{i}[t+1]=w_{ii}[t]x_{i}[t]+\sum_{j\in\mathcal{N}_{i}}w_{ij}[t]x_{j}[t],$ (1)
where $w_{ij}[t]$ is the weight assigned to agent $j$’s value by agent $i$, and where the following conditions are assumed for $w_{ij}[t]$ for $\forall i\in\mathcal{V}$ and $t\in\mathbb{Z}_{\geq 0}$:
  • $w_{ij}[t]=0$ $\forall j\notin\mathcal{N}_{i}\cup\{i\}$,
  • $w_{ij}[t]\geq\alpha$, $10\leq\alpha<$ 1 $\forall j\in\mathcal{N}_{i}\cup\{i\}$,
  • $\sum_{j=1}^{n}w_{ij}[t]=1$
Through the protocol given by (1), agents’ states reach consensus as long as the graph has a rooted-out branching (i.e. there exists a node that has paths to all other nodes in the graph) [38, 39]. However, as discussed in [4], this algorithm cannot achieve consensus in the presence of misbehaving agents. In our paper, we focus on two types of misbehaving agents, whose definitions are given below:
Definition 1 (malicious agent)
An agent $i\in\mathcal{V}$ is called malicious if it does not follow the nominal update protocol (1) at some time step $t$.
Definition 2 (Byzantine agent)
An agent $i\in\mathcal{V}$ is called Byzantine if it does not follow the nominal update protocol (1), or if it does not send the same value to all of its neighbors at some time step $t$.
Agents that are not misbehaving are called normal agents. There are numerous scope models to describe the number of misbehaving agents in a network [4]. The misbehaving agents are assumed to adopt one model, and we discuss two of them below:
Definition 3 ($\mathbf{F}$-total)
A set $\mathcal{S}\subset\mathcal{V}$ is $\mathbf{F}$-total if it contains at most $F$ nodes in the graph (i.e. $|\mathcal{S}|\leq F)$.
Definition 4 ($\mathbf{F}$-local)
A set $\mathcal{S}\subset\mathcal{V}$ is $\mathbf{F}$-local if all other nodes have at most $F$ nodes of $\mathcal{S}$ as their neighbors (i.e. $|\mathcal{N}_{i}\cap\mathcal{S}|\leq F$, $\forall i\in\mathcal{V}\setminus\mathcal{S}$).
In response to various threat and scope models, algorithms on resilient consensus [4, 15, 40, 5, 12, 13] have become very popular. In particular, the W-MSR (Weighted-Mean Subsequent Reduced) algorithm [4] with the parameter $F$ guarantees normal agents to achieve asymptotic consensus with $F$-total or $F$-local misbehaving agents under certain assumed topological properties of the communication graph. We review these properties and other relevant concepts in the next section.
2.3 $r$- and $(r,s)$-Robustness
Definition 5 ($\mathbf{r}$-reachable)
Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be a graph and $\mathcal{S}$ be a nonempty subset of $\mathcal{V}$. The subset $\mathcal{S}$ is $\mathbf{r}$-reachable if $\exists i\in\mathcal{S}$ such that $|\mathcal{N}_{i}\backslash\mathcal{S}|\geq r$, where $r\in\mathbb{Z}_{>0}$.
Definition 6 ($\mathbf{r}$-robust)
A graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ is $\mathbf{r}$-robust if $\forall\mathcal{S}_{1},\mathcal{S}_{2}\subset\mathcal{V}$ where $\mathcal{S}_{1}\cap S_{2}=\emptyset$ and $\mathcal{S}_{1},\mathcal{S}_{2}\neq\emptyset$, at least one of them is $r$-reachable.
Definition 7 ($\mathbf{(r,s)}$-reachable)
Given a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ and a nonempty subset $\mathcal{S}\subset\mathcal{V}$, $\mathcal{S}$ is $\mathbf{(r,s)}$-reachable if there are at least $s\in\mathbb{Z}_{\geq 0}$ nodes in $\mathcal{S}$ with at least $r\in\mathbb{Z}_{>0}$ neighbors outside of $\mathcal{S}$; i.e., given $\mathcal{X}^{r}_{\mathcal{S}}=\{i\in\mathcal{S}:|\mathcal{N}_{i}\setminus\mathcal{S}|\geq r\}$, then $|\mathcal{X}^{r}_{\mathcal{S}}|\geq s$.
Definition 8 ($\mathbf{(r,s)}$-robust)
A graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ is $\mathbf{(r,s)}$-robust if for every pair of nonempty, disjoint subsets $\mathcal{S}_{1},\mathcal{S}_{2}\subset\mathcal{V}$, at least one of the following holds (recall $\mathcal{X}^{r}_{\mathcal{S}_{k}}=\{i\in\mathcal{S}_{k}:|\mathcal{N}_{i}\setminus\mathcal{S}_{k}|\geq r\}$ for $k\in\{1,2\}$):
  • $|\mathcal{X}^{r}_{\mathcal{S}_{1}}|=|\mathcal{S}_{1}|$;
  • $|\mathcal{X}^{r}_{\mathcal{S}_{2}}|=|\mathcal{S}_{2}|$;
  • $|\mathcal{X}^{r}_{\mathcal{S}_{1}}|$ + $|\mathcal{X}^{r}_{\mathcal{S}_{2}}|\geq s$.
It can be stated that $(r,s)$-robustness is a generalization of $r$-robustness. If a graph is $(r,s)$-robust, it implies it is also $r$-robust. Conversely, if a graph is $r$-robust, it implies it is at least $(r,1)$-robust. These notions of robustness quantify how many misbehaving agents the network can tolerate while ensuring normal agents in a network to reach consensus on a value within the convex hull of their initial values through the W-MSR algorithm. Specifically, $(2F+1)$-robustness provides a sufficient condition for normal agents to achieve consensus in the presence of $F$-total and $F$-local malicious and Byzantine agents. Similarly, $(F+1,F+1)$-robustness serves as a necessary and sufficient condition for tolerating $F$-total malicious agents [4].
2.4 Problem Statement
In this paper, we aim to construct graphs with a given number of nodes that achieve the maximum $r$- and $(r,s)$- robustness while minimizing the number of edges. Consider a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ where $|\mathcal{V}|=n$. Before formally defining our problems, we first define two classes of graphs below:
Definition 9 ($\mathbf{r}$-MERG)
A graph $\mathcal{G}$ is $\mathbf{r}$-Minimal Edge Robust Graph, or $\mathbf{r}$-MERG if $\mathcal{G}$ achieves $r$-robustness with the fewest edges possible.
Definition 10 ($\mathbf{(r,s)}$-MERG)
A graph $\mathcal{G}$ is $\mathbf{(r,s)}$-Minimal Edge Robust Graph, or $\mathbf{(r,s)}$-MERG if $\mathcal{G}$ achieves $(r,s)$-robustness with the fewest edges possible.
The $r$-MERGs and $(r,s)$-MERGs are general classes of graphs that achieve certain robustness levels with minimal sets of edges. Note that in [37] we have provided similar definitions but due to the conflict of names, we are redefining with different names. In our paper, we are interested in finding the special classes of these, i.e., those that achieve maximum robustness with the fewest edges. Maximum $r$-robustness $\mathcal{G}$ can achieve is $r=\lceil\frac{n}{2}\rceil$ [4]. Using this fact, we now formally introduce two problems of our interest:
Problem 1
Given a desired number of nodes $n\in\mathbb{Z}_{>0}$, construct $\lceil\frac{n}{2}\rceil$-MERGs.
Problem 2
Given a desired number of nodes $n\in\mathbb{Z}_{>0}$, construct $(\lceil\frac{n}{2}\rceil,\lceil\frac{n}{2}\rceil)$-MERGs.
Remark 1
While the actual maximum $(r,s)$-robustness $\mathcal{G}$ can achieve is $(\lceil\frac{n}{2}\rceil,n)$, this level can only be achieved by complete graphs [4, Lemma 4], which are rather trivial. Instead, we focus our analysis on the case where $r=s=\lceil\frac{n}{2}\rceil$, as $(\lceil\frac{n}{2}\rceil,\lceil\frac{n}{2}\rceil)$-robust network can tolerate up to $\lceil\frac{n}{2}\rceil-1$ malicious agents, which is the maximum number of malicious agents the network with $n$ nodes can tolerate through the W-MSR in practice.
For ease of notation, we define $\gamma=\lceil\frac{n}{2}\rceil$ for the remainder of the paper. It is important to note that analyzing graph structures with maximum robustness requires distinct analyses for even and odd values of $n$. In fact, we can describe $n$ in terms of $\gamma$ as follows: $n=2\gamma-1$ for odd $n$, and $n=2\gamma$ for even $n$. Note that for a fixed $\gamma$, the even case yields an $n$ value one greater than the odd case. Despite this distinction, the solution procedures remain the same. Our analysis takes place in two main steps: First, we establish the lower bounds on the edge counts for a graph with $n$ nodes to achieve maximum robustness ($\gamma$- and $(\gamma,\gamma)$-robustness). Second, we leverage these lower bounds to develop construction mechanisms that are proven to construct $\gamma$- and $(\gamma,\gamma)$-MERGs.
3 Bounds on Edge Counts for $\gamma$-MERGs
This section addresses the first step toward constructing $\gamma$-MERGs by examining the minimum number of edges required for a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ with $|\mathcal{V}|=n$ to achieve $\gamma=\lceil\frac{n}{2}\rceil$-robustness. Due to the ceiling function involving $\gamma$, it is necessary to separately analyze cases where $n$ is odd or even. Thus, this section is divided into two cases. For the first case, we find the lower bounds on edge counts for $\gamma$-robust graphs when $n$ is odd. Then, for the second case, we do the same for $\gamma$-robust graphs when $n$ is even. These are formally established in Lemma 3.2 and Lemma 3.10, respectively. These two lemmas will be later used in Section 4 to construct $\gamma$-MERGs with any $n$.
Before analyzing the lower bounds on the edge counts for $\gamma$-robust graphs, we first introduce the concept of clique:
Definition 11 (clique [41])
A clique $\mathcal{C}$ of a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ is a subset $\mathcal{C}\subseteq\mathcal{V}$ such that $\forall i,j\in\mathcal{C}$ with $i\neq j$, we have $(i,j)\in\mathcal{E}$.
In general, $k$-clique refers to a clique consisting of $k$ nodes. The clique with the largest cardinality in $\mathcal{G}$ is called the maximum clique. Cliques are the building blocks of our analysis in this section, as we examine the necessary structures of $r$-robust graphs in general graph settings. We will use cliques to discover some necessary graph structures within $\gamma$-robust graphs, which will help us establish the lower bounds on edge counts.
3.1 Case 1: Graphs with Odd $n$
We now examine the minimum required size of the largest clique that any $\gamma$-robust graph with an odd number of nodes $n$ must contain in order to satisfy the conditions of $r$-robustness (Definition 6).
Figure 1: These figures illustrate the first two iterations in the proof of Lemma 1. Figure (a) shows the initial configuration, where node $v_{1}\in\mathcal{V}$ has $\gamma$ neighbors (colored in yellow). Figure (b) shows the first iteration where node $v_{2}\in\mathcal{S}_{1}$ is connected to $\gamma-1$ purple nodes and to $v_{1}$ in $\mathcal{S}_{2}$. The green nodes $v_{1}$ and $v_{2}$ form a $2$-clique. Note we maintain $|\mathcal{S}_{1}|=\gamma-1$ and $|\mathcal{S}_{2}|=\gamma$ from figures (b) to (e). Figure (c) shows $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ after the completion of the first step, where $\gamma-1$ nodes in $\mathcal{S}_{1}$ (including $v_{2}$) have been swapped with $\gamma-1$ purple nodes in $\mathcal{S}_{2}\setminus\{v_{1}\}$. Then, the start of the second step is shown in Figure (d), where another node $v_{3}\in\mathcal{S}_{1}$ has edges with $\gamma-2$ red nodes and with $v_{1},v_{2}$ in $\mathcal{S}_{2}$. Here, the green nodes $v_{1},v_{2},$ and $v_{3}$ form a $3$-clique. Figure (e) shows $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ after $\gamma-2$ nodes in $\mathcal{S}_{1}$ (including $v_{3}$) have been swapped with $\gamma-2$ red nodes in $\mathcal{S}_{2}\setminus\{v_{1},v_{2}\}$. This process continues until $(\gamma+1)$-clique is formed.
Lemma 1
Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be a graph of $n$ nodes where $n$ is odd, and let $\gamma=\lceil\frac{n}{2}\rceil$. Then, if $\mathcal{G}$ is $\gamma$-robust, it must contain a $(\gamma+1)$-clique.
Proof 3.1.
In this proof, we will show our argument by constructing two subsets $\mathcal{S}_{1},\mathcal{S}_{2}\subset\mathcal{V}$ such that $\mathcal{G}$ contains a $(\gamma+1)$-clique. Since $\mathcal{G}$ is $\gamma$-robust, it holds that for any pair of disjoint, non-empty subsets $\mathcal{S}_{1}$, $\mathcal{S}_{2}$ of $\mathcal{V}$, at least one of them is $\gamma$-reachable. Hence, by Definition 6, for any node in $\mathcal{V}=\{v_{1},\dots,v_{n}\}$, $|\mathcal{N}_{v_{i}}|\geq\gamma$ must hold. WLOG, let $v_{1}\in\mathcal{V}$ be a node, as shown on Figure 1 (a). Then because $n=2\gamma-1$, we can construct $\mathcal{S}_{1}$ to be the set of $\gamma-1$ neighbors of $v_{1}$ and $\mathcal{S}_{2}$ to be the set of the remaining $\gamma$ nodes including $v_{1}$ such that $|\mathcal{S}_{1}|=\gamma-1$, $|\mathcal{S}_{2}|=\gamma$, and $\mathcal{S}_{1}\cap\mathcal{S}_{2}=\emptyset$. This enforces $\mathcal{S}_{1}$ to be $\gamma$-reachable.
Since $\mathcal{S}_{1}$ is $\gamma$-reachable, $\exists v_{2}\in\mathcal{S}_{1}$ such that $|\mathcal{N}_{v_{2}}\cap\mathcal{S}_{2}|=|\mathcal{N}_{v_{2}}\setminus\mathcal{S}_{1}|\geq\gamma$. Then in the first step, $v_{2}$ has an edge with $v_{1}\in\mathcal{S}_{2}$, forming a $2$-clique as shown on Figure 1 (b). We then swap $\gamma-1$ nodes in $\mathcal{S}_{1}$ including $v_{2}$ with $\gamma-1$ nodes in $\mathcal{S}_{2}\setminus\{v_{1}\}$ (colored in purple in Figure 1). Then in the second step, since $\mathcal{S}_{1}$ is $\gamma$-reachable even after $v_{2}$ is swapped out, $\exists v_{3}\in\mathcal{S}_{1}$ such that $|\mathcal{N}_{v_{3}}\cap\mathcal{S}_{2}|\geq\gamma$. By construction, since $v_{1}$ and $v_{2}$ are among the $\gamma$ nodes in $\mathcal{S}_{2}$, $v_{3}$ forms a $3$-clique with $v_{1}$ and $v_{2}$, as shown on Figure 1 (d). We then swap $\gamma-2$ nodes in $\mathcal{S}_{1}$ including $v_{3}$ with $\gamma-2$ nodes in $\mathcal{S}_{2}\setminus\{v_{1},v_{2}\}$ (colored in red in Figure 1). Again, in the third step, since $\mathcal{S}_{1}$ is $\gamma$-reachable even after $v_{3}$ is swapped out, $\exists v_{4}\in\mathcal{S}_{1}$ that has edges with all $\gamma$ nodes in $\mathcal{S}_{2}$. Since $v_{1},v_{2},v_{3}\in\mathcal{S}_{2}$ are three of the $\gamma$ nodes to have edges with $v_{4}$, the nodes $v_{1},v_{2},v_{3},v_{4}$ form a $4$-clique. We then swap $\gamma-3$ nodes in $\mathcal{S}_{1}$ including $v_{4}$ with $\gamma-3$ nodes in $\mathcal{S}_{2}\setminus\{v_{1},v_{2},v_{3}\}$, and thus the construction continues.
Likewise, in the $k^{\text{th}}$ step, $\exists v_{k+1}\in\mathcal{S}_{1}$ such that $|\mathcal{N}_{v_{k+1}}\cap\mathcal{S}_{2}|\geq\gamma$ and thus forms a $(k+1)$-clique with $v_{1},v_{2},\cdots,v_{k-1},v_{k}\in\mathcal{S}_{2}$. We then swap $\gamma-k$ nodes in $\mathcal{S}_{1}$ including $v_{k+1}$ with $\gamma-k$ nodes in $\mathcal{S}_{2}\setminus\{v_{1},\cdots,v_{k}\}$. At the end of $\gamma-1^{\text{th}}$ step, $\mathcal{S}_{2}$ contains $v_{1},\cdots,v_{\gamma}$ that form a $\gamma$-clique. Finally, since $\mathcal{S}_{1}$ is $\gamma$-reachable, it contains a node $v_{\gamma+1}$ that has edges with all nodes $v_{1},\cdots,v_{\gamma}\in\mathcal{S}_{2}$. Then, $\{v_{\gamma+1}\}\cup\mathcal{S}_{2}$ forms a ($\gamma+1)$-clique.
Lemma 1 states that for a graph with odd $n$ to achieve $\gamma$-robustness (maximum $r$-robustness), it must contain a clique size of $\gamma+1$ or more nodes. This effectively builds a necessary subgraph structure for further analysis. Note that the existence of $(\gamma+1)$-clique implies that $\mathcal{G}$ must have as many edges as to construct a $(\gamma+1)$-clique, i.e., $|\mathcal{E}|\geq\frac{\gamma(\gamma+1)}{2}$. Stemming from this reasoning, we finalize the lower bound on edge counts for the case of odd $n$:
Lemma 3.2.
Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be a graph of $n$ nodes where $n$ is odd, and let $\gamma=\lceil\frac{n}{2}\rceil$. Then, if $\mathcal{G}$ is $\gamma$-robust,
$|\mathcal{E}|\geq\frac{3\gamma(\gamma-1)}{2}.$ (2)
Proof 3.3.
Since $\mathcal{G}$ is $\gamma$-robust, it holds that for any pair of disjoint, non-empty subsets $\mathcal{S}_{1}$, $\mathcal{S}_{2}\subset\mathcal{V}$, at least one of them is $\gamma$-reachable (by Definition 6). From Lemma 1, we know that $\mathcal{V}$ must contain a $(\gamma+1)$-clique $\mathcal{C}$. Let $\mathcal{C}^{\prime}=\mathcal{V}\setminus\mathcal{C}$. This then implies $|\mathcal{C}|=\gamma+1$ and $|\mathcal{C}^{\prime}|=\gamma-2$.
WLOG, let $\mathcal{S}_{1}=\mathcal{C}^{\prime}$ and $\mathcal{S}_{2}=\mathcal{C}$. Then, $|\mathcal{S}_{1}|=\gamma-2$ and $|\mathcal{S}_{2}|=\gamma+1$. This implies $\mathcal{S}_{1}$ has to be $\gamma$-reachable. In other words, there must exist $v_{1}\in\mathcal{S}_{1}$ such that $|\mathcal{N}_{v_{1}}\cap\mathcal{S}_{2}|=|\mathcal{N}_{v_{1}}\setminus\mathcal{S}_{1}|\geq\gamma$. Therefore, for such $v_{1}$ to exist, we need $\gamma$ edges that connect $v_{1}$ with $\gamma$ nodes in $\mathcal{S}_{2}$: $|\mathcal{E}|\geq\frac{\gamma(\gamma+1)}{2}+\gamma$. Note that $v_{1}\in\mathcal{S}_{1}$ is the node that makes $\mathcal{S}_{1}$ $\gamma$-reachable. Therefore, moving $v_{1}$ from $\mathcal{S}_{1}$ to $\mathcal{S}_{2}$ requires the new $\mathcal{S}_{1}$, which is now without $v_{1}$, to contain another node $v_{2}$ that has at least $\gamma$ neighbors in $\mathcal{S}_{2}$. This again requires at least $\gamma$ additional edges (i.e. $|\mathcal{E}|\geq\frac{\gamma(\gamma+1)}{2}+2\gamma$). We then move $v_{2}$ from $\mathcal{S}_{1}$ to $\mathcal{S}_{2}$, which necessitates $\mathcal{S}_{1}$ to have another node $v_{3}$ to have $\gamma$ neighbors in $\mathcal{S}_{2}$.
We continue this process of (i) connecting a node $v_{i}\in\mathcal{S}_{1}$ with any $\gamma$ nodes in $\mathcal{S}_{2}$ and (ii) moving $v_{i}$ from $\mathcal{S}_{1}$ to $\mathcal{S}_{2}$ in the $i^{\text{th}}$ step until $\mathcal{S}_{1}$ becomes empty. Then, we can repeat these two steps $\gamma-2$ times, as we initially had $|\mathcal{S}_{1}|=\gamma-2$. In other words, for each $\gamma-2$ pair of $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$, we require $\gamma$ additional edges. Since we started from the setting where $\mathcal{S}_{2}=\mathcal{C}$ and forced $|\mathcal{E}|$ to increase at least a total of $\gamma(\gamma-2)$, $|\mathcal{E}|\geq\frac{\gamma(\gamma+1)+2\gamma(\gamma-2)}{2}\Rightarrow|\mathcal{E}|\geq\frac{3\gamma(\gamma-1)}{2}$.
From Lemma 3.2, we can conclude that any $\gamma$-robust graphs with odd $n$, i.e., graphs with an odd number of nodes that achieve the maximum $r$-robustness, must have $\frac{3\gamma(\gamma-1)}{2}$ or more edges. In fact, Theorem 4.16 later shows that the bound in the lemma is tight and cannot be further improved. Furthermore, using Lemma 3.2, we can derive a useful corollary:
Corollary 3.4.
Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be an $r$-robust graph, where $r\in\mathbb{Z}_{>0}$. Then,
$|\mathcal{E}|\geq\frac{3r(r-1)}{2}.$ (3)
Proof 3.5.
In general, the more nodes a graph has, the more edges it must have to maintain a given robustness. From [30], it is shown that $12r-$ 1 is the smallest number of nodes a graph must have to be $r$-robust. In other words, $r$-robustness is the maximum robustness a graph with $12r-$ 1 nodes can achieve. Thus, Lemma 3.2 can be restated: any $r$-robust graph with $12r-$ 1 nodes must have at least $\frac{3r(r-1)}{2}$ edges. Consequently, any graph $\mathcal{G}$ with $12r-$ 1 or more nodes must have at least $\frac{3r(r-1)}{2}$ edges to be $r$-robust.
Remark 3.6.
Corollary 3.4 reveals that the bound given in Lemma 3.2 can be applied beyond $\gamma$-robustness. The bound in fact serves as a necessary condition for any $r$-robust graphs of any number of nodes $n$, ensuring that graphs with edges less than $\frac{3r(r-1)}{2}$ cannot be $r$-robust. One useful implication of this result is that it provides a quick and easy way to upper-bound a graph’s robustness level based solely on its number of edges. Computing $r$-robustness of a graph is in general a very difficult problem [8, 21, 42]. While various efficient methods have been proposed to approximate robustness levels [22, 23, 24], these approaches would still involve significant computational effort as graph sizes increase and frequently lack formal guarantees. In contrast, verifying whether a given graph satisfies the bound (2) is computationally efficient. It can also accelerate existing methods by reducing their search space through the upper bound on robustness provided by this result beforehand.
3.2 Case 2: Graphs with Even $n$
Now we continue our analysis on the edge counts for the graphs with even $n$. Since $n$ is even, $\gamma=\frac{n}{2}$. This implies that, for the same value of $\gamma$, the value of $n$ is one greater when $n$ is even compared to when it is odd. Because there is one more node, the analysis involves more steps. However, the overall approach is similar to the previous case: we will first determine the minimum size of the maximum clique that $\mathcal{G}$ must contain to be $\gamma$-robust.
Lemma 3.7.
Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be a graph with $n$ nodes where $n$ is even, and let $\gamma=\frac{n}{2}$. Then, if $\mathcal{G}$ is $\gamma$-robust, it must contain a $\big{(}\lfloor\frac{\gamma+4}{2}\rfloor\big{)}$-clique.
Due to length, the proof for Lemma 3.7 is given in the Appendix (Section 9). Comparing Lemma 1 and 3.7, it can be seen that the size of the necessary maximum clique in a $\gamma$-robust graph we can guarantee decreases as its number of node increases. This is reasonable, as the number of possible graph configurations increases exponentially with the number of nodes, thereby making the analysis more complex. Due to the increased complexity, unlike the analysis in the previous subsection for the odd $n$ case, the study of $\gamma$-robust graphs with even $n$ requires topological structures beyond the maximum clique. The following lemma presents the additional necessary structures.
Lemma 3.8.
Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be a graph of $n$ nodes where $n$ is even, and let $\gamma=\frac{n}{2}$. Then, if $\mathcal{G}$ is $\gamma$-robust, $\mathcal{G}$ contains an induced subgraph $\mathcal{G}_{S}=(\mathcal{V}_{S},\mathcal{E}_{S})$ such that $|\mathcal{V}_{S}|=\gamma+1$ and $|\mathcal{E}_{S}|\geq\lfloor{\frac{\gamma^{2}+2}{2}}\rfloor$.
Proof 3.9.
Since $\mathcal{G}$ is $\gamma$-robust, for any pair of disjoint, non-empty subsets $\mathcal{S}_{1}$, $\mathcal{S}_{2}\subset\mathcal{V}$, at least one of them is $\gamma$-reachable. From Lemma 3.7, $\mathcal{G}$ contains a $\big{(}\lfloor\frac{\gamma+4}{2}\rfloor\big{)}$-clique $\mathcal{C}$. WLOG, let $\mathcal{S}_{2}$ contain $\gamma+1$ nodes including $\mathcal{C}$, and let $\mathcal{S}_{1}$ contain the remaining $\gamma-1$ nodes. Initially, we set $\mathcal{V}_{S}=\mathcal{C}$. Then, $|\mathcal{E}_{S}|=\big{(}\frac{1}{2}\big{)}\big{(}\lfloor\frac{\gamma+4}{2}\rfloor\big{)}\big{(}\lfloor\frac{\gamma+2}{2}\rfloor\big{)}$. Note there are total $n=2\gamma$ nodes. Since $|\mathcal{S}_{1}|=\gamma-1$ and $|\mathcal{S}_{2}|=\gamma+1$, $\mathcal{S}_{1}$ must be $\gamma$-reachable.
By Lemma 3.7, we have a $2$-clique and $3$-clique when $\gamma=1$ and $\gamma=2$ respectively, satisfying the statement in the lemma for $\gamma=1,2$.
Now, we consider $\gamma\geq 3$. With this setup, since $\mathcal{S}_{1}$ is $\gamma$-reachable, $\exists v_{1}\in\mathcal{S}_{1}$ such that $|\mathcal{N}_{v_{1}}\cap\mathcal{S}_{2}|=|\mathcal{N}_{v_{1}}\setminus\mathcal{S}_{1}|\geq\gamma$. We swap $v_{1}\in\mathcal{S}_{1}$ with any node in $\mathcal{S}_{2}\setminus\mathcal{V}_{S}$. Note that $|\mathcal{V}_{S}|=\lfloor\frac{\gamma+4}{2}\rfloor$ at this point. Now, we add $v_{1}$ as a vertex of $\mathcal{G}_{S}$ (i.e. $\mathcal{V}_{S}$ now contains $v_{1}$). Then, $|\mathcal{E}_{S}|$ increases at least by $\lfloor\frac{\gamma+2}{2}\rfloor$, since $v_{1}$ may have no edge with one of the nodes in $\mathcal{V}_{S}$. Since $\mathcal{S}_{1}$ is $\gamma$-reachable even without $v_{1}$, $\exists v_{2}\in\mathcal{S}_{1}$ such that $|\mathcal{N}_{v_{2}}\cap\mathcal{S}_{2}|\geq\gamma$. Note that $|\mathcal{V}_{S}|=\lfloor\frac{\gamma+6}{2}\rfloor$ at this point. We swap $v_{2}\in\mathcal{S}_{1}$ with a node in $\mathcal{S}_{2}\setminus\mathcal{V}_{S}$ and add $v_{2}$ as a new vertex of $\mathcal{G}_{S}$ (i.e. $\mathcal{V}_{S}$ now contains $v_{2}$). This increases $|\mathcal{E}_{S}|$ at least by $\lfloor\frac{\gamma+4}{2}\rfloor$, as $v_{2}$ may have no edge with one of the nodes in $\mathcal{V}_{S}$. Then, since $\mathcal{S}_{1}$ is still $\gamma$-reachable, $\exists v_{3}\in\mathcal{S}_{1}$ such that $|\mathcal{N}_{v_{3}}\cap\mathcal{S}_{2}|\geq\gamma$. Note that $|\mathcal{V}_{S}|=\lfloor\frac{\gamma+8}{2}\rfloor$. Again, we swap $v_{3}\in\mathcal{S}_{1}$ with a node in $\mathcal{S}_{2}\setminus\mathcal{V}_{S}$ and add $v_{3}$ as a new vertex of $\mathcal{G}_{S}$. Then, $|\mathcal{E}_{S}|$ increases at least by $\lfloor\frac{\gamma+6}{2}\rfloor$.
This process continues until when $v_{p}\in\mathcal{S}_{1}$, $p=\gamma+1-\lfloor\frac{\gamma+4}{2}\rfloor$, that makes $\mathcal{S}_{1}$ $\gamma$-reachable and gets swapped with a node in $\mathcal{S}_{2}\setminus\mathcal{V}_{S}$ where $\mathcal{V}_{S}=\mathcal{C}\cup\{v_{1},v_{2},\cdots,v_{p-1}\}$. Here, adding $v_{p}$ as a new vertex of $\mathcal{G}_{S}$ increases $|\mathcal{E}_{S}|$ at least by $\gamma-1$. At this point, $\mathcal{S}_{2}=\mathcal{V}_{S}$ and $|\mathcal{V}_{S}|=\gamma+1$. Then, $|\mathcal{E}_{S}|\geq\big{(}\frac{1}{2}\big{)}\big{(}\lfloor\frac{\gamma+4}{2}\rfloor\big{)}\big{(}\lfloor\frac{\gamma+2}{2}\rfloor\big{)}+(\lfloor\frac{\gamma+2}{2}\rfloor)+(\lfloor\frac{\gamma+4}{2}\rfloor)+(\lfloor\frac{\gamma+6}{2}\rfloor)+\cdots+(\gamma-2)+(\gamma-1)$.
We remove the floor functions: if $\gamma$ is even, $|\mathcal{E}_{S}|\geq\big{(}\frac{1}{2}\big{)}\big{(}\frac{\gamma+4}{2}\big{)}\big{(}\frac{\gamma+2}{2}\big{)}+\big{(}\frac{\gamma+2}{2}\big{)}+\big{(}\frac{\gamma+4}{2}\big{)}+\cdots+(\gamma-2)+(\gamma-1)=\frac{\gamma^{2}+2}{2}$. If $\gamma$ is odd, $|\mathcal{E}_{S}|\geq\big{(}\frac{1}{2}\big{)}\big{(}\frac{\gamma+3}{2}\big{)}\big{(}\frac{\gamma+1}{2}\big{)}+\big{(}\frac{\gamma+1}{2}\big{)}+\big{(}\frac{\gamma+3}{2}\big{)}+\cdots+(\gamma-2)+(\gamma-1)=\frac{\gamma^{2}+1}{2}$. Thus, we have $|\mathcal{V}_{S}|=\gamma+1$ and $|\mathcal{E}_{S}|\geq\lfloor{\frac{\gamma^{2}+2}{2}}\rfloor$.
Lemma 3.8 effectively provides a necessary induced subgraph structure that $\mathcal{G}$ with even $n$ must have to be $\gamma$-robust. Now, using this lemma, we present the following:
Lemma 3.10.
Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be a graph of $n$ nodes where $n$ is even, and let $\gamma=\frac{n}{2}$. Then, if $\mathcal{G}$ is $\gamma$-robust,
$|\mathcal{E}|\geq\left\lfloor{\frac{\gamma(3\gamma-2)+2}{2}}\right\rfloor.$ (4)
Proof 3.11.
The approach of the proof is identical to that of Lemma 3.2, except for the fact that in this proof, we are starting with an induced subgraph $\mathcal{G}_{S}=(\mathcal{V}_{S},\mathcal{E}_{S})$ such that $|\mathcal{V}_{S}|=\gamma+1$ and $|\mathcal{E}_{S}|\geq\lfloor{\frac{r^{2}+2}{2}}\rfloor$ (which comes from Lemma 3.8).
Let $\mathcal{S}_{2}=\mathcal{V}_{S}$ and $\mathcal{S}_{1}=\mathcal{V}\setminus\mathcal{V}_{S}$, which means $|\mathcal{S}_{1}|=\gamma-1$ and $|\mathcal{S}_{2}|=\gamma+1$. This enforces $\mathcal{S}_{1}$ to be $\gamma$-reachable. Then, $\exists v_{1}\in\mathcal{S}_{1}$ such that $|\mathcal{N}_{v_{1}}\cap\mathcal{S}_{2}|=|\mathcal{N}_{v_{1}}\setminus\mathcal{S}_{1}|\geq\gamma$. This requires at least $\gamma$ edges (i.e. $|\mathcal{E}|\geq|\mathcal{E}_{S}|+\gamma$). We move $v_{1}$ from $\mathcal{S}_{1}$ to $\mathcal{S}_{2}$. Then since $\mathcal{S}_{1}$ is $\gamma$-reachable even without $v_{1}$, $\exists v_{2}\in\mathcal{S}_{1}$ such that $|\mathcal{N}_{v_{2}}\cap\mathcal{S}_{2}|\geq\gamma$. This requires at least $\gamma$ new edges (i.e. $|\mathcal{E}|\geq|\mathcal{E}_{S}|+2\gamma$).
We can continue this process of (i) drawing edges between $v_{i}\in\mathcal{S}_{1}$ and $\gamma$ nodes in $\mathcal{S}_{2}$ and (ii) moving $v_{i}\in\mathcal{S}_{1}$ into $\mathcal{S}_{2}$ until $\mathcal{S}_{1}$ becomes empty. Then, we can do these steps $\gamma-1$ times, as we initially have $|\mathcal{S}_{1}|=\gamma-1$. In other words, for each $\gamma-1$ pair of $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$, we require $\gamma$ additional edges. Since we have started from the setting where $\mathcal{S}_{2}=\mathcal{V}_{S}$, we have $|\mathcal{E}|\geq|\mathcal{E}_{S}|+\gamma(\gamma-1)$. Therefore, if $\gamma$ is even, $|\mathcal{E}|\geq\frac{\gamma^{2}+2}{2}+\gamma(\gamma-1)=\frac{\gamma(3\gamma-2)+2}{2}$. If $\gamma$ is odd, and $|\mathcal{E}|\geq\frac{\gamma^{2}+1}{2}+\gamma(\gamma-1)=\frac{\gamma(3\gamma-2)+1}{2}$.
Remark 3.12.
While the primary goal of this section has been to establish bounds on the edge counts for $\gamma$-robust graphs, we also have identified some interesting graph structures that are necessary for $\gamma$-robust graphs. These structures arise solely from the definition of $r$-robustness (Definition 6) and provide topological representations of what $r$-robustness requires. A more systematic exploration of these necessary structures could yield valuable insights, particularly into how to relax the often dense connectivity requirements associated with robustness. Investigating this direction further presents a promising avenue for future work.
From Lemma 3.10, we can obtain a corollary similar to Corollary 3.4:
Corollary 3.13.
Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be an $r$-robust graph, where $r\in\mathbb{Z}_{>0}$ and $|\mathcal{V}|=n$ is even. Then,
$|\mathcal{E}|\geq\left\lfloor{\frac{r(3r-2)+2}{2}}\right\rfloor.$ (5)
Proof 3.14.
In general, the more nodes a graph has, the more edges it must have to maintain a certain robustness. From [30], it can be seen shown that $2r$ is the smallest even number of nodes a graph to have to be $r$-robust. In other words, $r$-robustness is the maximum robustness a graph with $2r$ nodes can achieve. Then by Lemma 3.10, a graph of even number of nodes needs at least $\left\lfloor{\frac{r(3r-2)+2}{2}}\right\rfloor$ edges to be $r$-robust.
Remark 3.15.
We compare the bounds established in Lemmas 3.2 and 3.10 to observe that the lower bound in the latter is strictly stronger. In other words, while the latter also applies to $\gamma$-robust graphs with even $n$, the former offers a tighter condition. This is also reflected in Corollary 3.4 and 3.13, where the latter provides a stronger necessary condition on the edge counts for any graphs with even $n$ to become $r$-robust than the former does.
4 Construction of $\gamma$-MERGs
In the previous section, we have analyzed and established the lower bounds on edge counts that graphs of $n$ nodes must satisfy in order to achieve $\gamma$-robustness (where $\gamma=\lceil\frac{n}{2}\rceil$ is the maximum robustness). Proceeding with the lower bounds, we construct $\gamma$-MERGs, or graphs that provably achieve $\gamma$-robustness with a minimal set of edges. This is formally given in Theorem 4.16, as our first main result:
Theorem 4.16.
Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be a graph with $n$ nodes, and let $\gamma=\lceil\frac{n}{2}\rceil$. Then, $\mathcal{G}$ is a $\gamma$-MERG under the following construction mechanisms:
  • If $n$ is odd, let $\mathcal{X}\subset\mathcal{V}$ be a set of $\gamma+1$ nodes that form a $(\gamma+1)$-clique. Connect each of the remaining $\gamma-1$ nodes to any $\gamma$ nodes in $\mathcal{X}$.
  • If $n$ is even, let $\mathcal{X}\subset\mathcal{V}$ be a set of $\gamma$ nodes, each of which is adjacent to all other nodes in $\mathcal{V}$. Select $\lceil\frac{\gamma-2}{2}\rceil$ disjoint pairs of nodes in $\mathcal{X}$ and remove the edge between each pair.
Proof 4.17.
To prove that $\mathcal{G}$ is $\gamma$-MERG, we need to prove two things for both odd and even values of $n$: (i) it is $\gamma$-robust and (ii) its edge count $|\mathcal{E}|$ equals to the lower bounds specified in Lemma 3.2 or Lemma 3.10, depending on the value of $n$.
Odd $\mathbf{n}$: We first prove that it is $\gamma$-robust. Let $\mathcal{X}^{\prime}=\mathcal{V}\setminus\mathcal{X}$. Then, $|\mathcal{X}|=\gamma+1$ and $|\mathcal{X}^{\prime}|=\gamma-2$. WLOG, let $\mathcal{S}_{1},\mathcal{S}_{2}\in\mathcal{V}$ be nonempty subsets such that $\mathcal{S}_{1}\cap\mathcal{S}_{2}=\emptyset$ and $|\mathcal{S}_{1}|\leq|\mathcal{S}_{2}|$. There are two cases.
1) If $\mathcal{X}\cap\mathcal{S}_{1}=\emptyset$, let a node $i\in\mathcal{X}^{\prime}\cap\mathcal{S}_{1}$. Since $\mathcal{X}\cap\mathcal{S}_{1}=\emptyset$ and $\mathcal{N}_{i}\subset\mathcal{X}\subseteq\mathcal{S}_{2}$ by construction, $|\mathcal{N}_{i}\setminus\mathcal{S}_{1}|\geq\gamma$, making $\mathcal{S}_{1}$ $\gamma$-reachable.
2) If $\mathcal{X}\cap\mathcal{S}_{1}\neq\emptyset$, denote $|\mathcal{X}\cap\mathcal{S}_{1}|=p\leq\gamma-1$. Then, we have $1\leq|\mathcal{X}^{\prime}\cap\mathcal{S}_{1}|\leq\gamma-1-p$, $|\mathcal{X}\setminus\mathcal{S}_{1}|=\gamma+1-p$, and $p-1\leq|\mathcal{X}^{\prime}\setminus\mathcal{S}_{1}|\leq\gamma-2$. For $i\in\mathcal{X}\cap\mathcal{S}_{1}$, $|(\mathcal{N}_{i}\setminus\mathcal{S}_{1})\cap\mathcal{X}|=|\mathcal{X}\setminus\{i\}|-(p-1)=\gamma-(p-1)$, as $\mathcal{X}$ forms a $(\gamma+1)$-clique. For $p=1$, $\exists i\in\mathcal{X}\cap\mathcal{S}_{1}$ such that $|\mathcal{N}_{i}\setminus\mathcal{S}_{1}|\geq|\mathcal{X}\setminus\{i\}|=\gamma$. For $p\geq 2$, note that by construction, every node in $\mathcal{X}^{\prime}$ has edges with any $\gamma$ nodes in $\mathcal{X}$, where $|\mathcal{X}|=\gamma+1$. Speaking differently, every node in $\mathcal{X}^{\prime}$ does not have an edge with only one node in $\mathcal{X}$. Thus, all nodes in $\mathcal{X}^{\prime}\setminus\mathcal{S}_{1}$ have edges with at least $p-1$ nodes in $\mathcal{X}\cap\mathcal{S}_{1}$, because it may not have an edge with only one node in $\mathcal{X}\cap\mathcal{S}_{1}$. Then, as $p-1\leq|\mathcal{X}^{\prime}\setminus\mathcal{S}_{1}|\leq\gamma-2$, there are at least a total of $(p-1)(p-1)$ edges between the nodes in $\mathcal{X}^{\prime}\setminus\mathcal{S}_{1}$ and $\mathcal{X}\cap\mathcal{S}_{1}$. Since $|\mathcal{X}\cap\mathcal{S}_{1}|=p$ and by Pigeonhole Principle, $\exists i\in\mathcal{X}\cap\mathcal{S}_{1}$ such that $|(\mathcal{N}_{i}\setminus\mathcal{S}_{1})\cap\mathcal{X}^{\prime}|\geq\big{\lceil}{\frac{(p-1)(p-1)}{p}}\big{\rceil}=\big{\lceil}{\frac{p^{2}-2p+1}{p}}\big{\rceil}=p-1$. Then, $\exists i\in\mathcal{X}\cap\mathcal{S}_{1}$ such that $|\mathcal{N}_{i}\setminus\mathcal{S}_{1}|=|(\mathcal{N}_{i}\setminus\mathcal{S}_{1})\cap\mathcal{X}|+|(\mathcal{N}_{i}\setminus\mathcal{S}_{1})\cap\mathcal{X}^{\prime}|\geq\gamma-(p-1)+p-1=\gamma$. Thus, whether $\mathcal{S}_{1}$ contains a node in $\mathcal{X}$ or not, $\mathcal{S}_{1}$ is always $\gamma$-reachable.
Now we show that $\mathcal{E}$ is a minimal set. $\mathcal{G}$ has $\frac{(\gamma+1)(\gamma)}{2}$ edges to form a $(\gamma+1)$-clique. Additionally, we introduce $\gamma(\gamma-2)$ edges for connecting each of the $\gamma-2$ nodes in $\mathcal{X}^{\prime}$ with $\gamma$ nodes in $\mathcal{X}$. Summing them up, we have $\frac{3\gamma(\gamma-1)}{2}$ which equals the lower bound given in Lemma 3.2.
Even $\mathbf{n}$: We first prove that it is $\gamma$-robust. Let $\mathcal{X}^{\prime}=\mathcal{V}\setminus\mathcal{X}$. Then, $|\mathcal{X}|=\gamma$ and $|\mathcal{X}^{\prime}|=\gamma$. WLOG, let $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ be nonempty subsets of $\mathcal{V}$ such that $\mathcal{S}_{1}\cap\mathcal{S}_{2}=\emptyset$ and $|\mathcal{S}_{1}|\leq|\mathcal{S}_{2}|$. That means $1\leq|\mathcal{S}_{1}|\leq\gamma$. There are two cases.
1) If $\mathcal{X}\cap\mathcal{S}_{1}=\emptyset$, let a node $i\in\mathcal{X}^{\prime}\cap\mathcal{S}_{1}$. Since $\mathcal{X}\cap\mathcal{S}_{1}=\emptyset$ and $\mathcal{N}_{i}=\mathcal{X}$, $|\mathcal{N}_{i}\setminus\mathcal{S}_{1}|\geq\gamma$, and thus $\mathcal{S}_{1}$ is $\gamma$-reachable.
2) If $\mathcal{X}\cap\mathcal{S}_{1}\neq\emptyset$, let a node $i\in\mathcal{X}\cap\mathcal{S}_{1}$. Note that after removing an edge among $\lceil\frac{\gamma-2}{2}\rceil$ pairs of nodes in $\mathcal{X}$, there exists at least one (two in case of even $\gamma$) node(s) that is connected to all other nodes in $\mathcal{G}$. In other words, $\exists v_{*}\in\mathcal{V}$ s.t. $|N_{v_{*}}|=2\gamma-1$. Now, when $11\leq|\mathcal{S}_{1}|\leq\gamma-$ 1, $|\mathcal{N}_{i}\setminus\mathcal{S}_{1}|\geq\gamma$, making $\mathcal{S}_{1}$ $\gamma$-reachable. In case $|\mathcal{S}_{1}|=|\mathcal{S}_{2}|=\gamma$, we know there exists $v_{*}$ in either $\mathcal{S}_{1}$ or $\mathcal{S}_{2}$ such that $|\mathcal{N}_{v_{*}}\setminus\mathcal{S}_{i}|\geq\gamma$, making $\mathcal{S}_{i}$ $\gamma$-reachable. Thus, whether $\mathcal{S}_{1}$ contains a node in $\mathcal{X}$ or not, either $\mathcal{S}_{1}$ or $\mathcal{S}_{2}$ is always $\gamma$-reachable.
Now, we examine its edge set. The $\gamma$ nodes in $\mathcal{X}$ connecting to every other node adds up to a total of $\frac{\gamma(\gamma-1)+2\gamma(\gamma)}{2}=\frac{3\gamma^{2}-\gamma}{2}$ edges. Then, if we subtract $\lceil\frac{\gamma-2}{2}\rceil$ from it, we get $\frac{\gamma(3\gamma-2)+2}{2}$ for even $\gamma$ and $\frac{\gamma(3\gamma-\gamma)+1}{2}$ for odd $\gamma$, which equals the lower bound given in Lemma 3.10.
Figure 2: Visualizations of $5$-MERGs with $9$ (left) and $10$ (right) nodes.
Theorem 4.16 provides a systematic procedure for constructing $\gamma$-MERGs, i.e., graphs that achieve maximum $r$-robustness with a minimal set of edges, giving a solution to Problem 1. As explicitly stated in the theorem, $\mathcal{G}$ contains an $(r+1)$-clique when $n$ is odd, further demonstrating Lemma 1. For even $n$, although not explicitly stated, we can observe that $\mathcal{G}$ contains a $\lfloor\frac{\gamma+4}{2}\rfloor$-clique, as given by Lemma 3.10. Furthermore, the theorem suggests that the bounds given in Lemma 3.2 and 3.10 are tight. In fact, one removal of any edge from a $\gamma$-MERG immediately drops the graph’s robustness level by at least one (as demonstrated in Sec. 7.2).
Some examples of $\gamma$-MERGs with $n=9$ (left) and $n=10$ (right) are visualized in Figure 2. The $5$-MERG with $n=9$ actually has a maximum clique of size $6$ (i.e., $\{0,1,2,3,4,5\}$), while the one with $n=10$ contains a $4$-clique as the maximum clique (i.e., $\{0,2,4,7\}$). These examples align with the necessary maximum clique sizes analyzed in Lemma 1 and 3.10: $(\gamma+1)$-cliques for odd $n$ and $\lfloor\frac{\gamma+4}{2}\rfloor$-cliques for even $n$.
5 Bounds on Edge Counts for $(\gamma,\gamma)$-MERGs
In Section 3 and 4, we have examined and constructed $\gamma$-MERGs to address Problem 1. In the following two sections, we focus on $(\gamma,\gamma)$-MERGs, graphs with a given number of nodes that achieve maximum $(r,s)$-robustness with a minimal set of edges, thereby addressing Problem 2.
Similar to what we have done in Section 3, since $\gamma=\lceil\frac{n}{2}\rceil$, we have two cases: $n=2\gamma-1$ for odd $n$ and $n=2\gamma$ for even $n$. However, $(\gamma,\gamma)$-MERGs for odd $n$ have already been characterized in [4, Lemma 4], which is given below for completeness:
Lemma 5.18.
[4] Let $n>1$ be odd and $\gamma=\lceil\frac{n}{2}\rceil$. Then, complete graphs are the only graphs with $n$ nodes that are $(\gamma,s)$-robust with $s\geq\lfloor\frac{n}{2}\rfloor$.
Lemma 5.18 essentially confirms that when $n$ is odd, the graph must be a complete graph to be a $(\gamma,\gamma)$-MERG. This also implies that when $n$ is odd, for a graph to be $(\gamma,\gamma)$-robust, it must have at least $\frac{(2\gamma-1)(2\gamma-2)}{2}=(\gamma-1)(2\gamma-1)$ edges.
5.1 Graphs with Even $n$
We now introduce the lower bounds on edge counts for $(\gamma,\gamma)$-robust graphs when $n$ is even. This is formally established in Lemma 5.22. This lemma will be later used in Section 6 to construct $(\gamma,\gamma)$-MERGs for any $n$. However, before that, we introduce another useful lemma from [4]:
Lemma 5.19.
Given an $(r,s)$-robust graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, with $0\leq r\leq\lceil\frac{n}{2}\rceil$ and $1\leq s\leq n$, the minimum degree of $\mathcal{G}$, $\delta_{\min}(\mathcal{G})$, is at least
$\delta_{\min}(\mathcal{G})\geq\begin{cases}2r-2&\text{if }s\geq r\\ r+s-1&\text{otherwise}\end{cases}.$ (6)
Because we consider $r=s=\gamma$ in our paper, what we really care in (6) is the inequality for the first case: $\delta_{\min}(\mathcal{G})\geq 2r-2$. Therefore, we have the following corollary:
Corollary 5.20.
Given an $(r,r)$-robust graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ of $n$ nodes, the minimum degree of $\mathcal{G}$, $\delta_{\min}(\mathcal{G})\geq 2(r-1)$. Furthermore, $|\mathcal{E}|\geq n(r-1)$.
Proof 5.21.
As the direct application of Lemma 6, we can conclude $\delta_{\min}(\mathcal{G})\geq 2(r-1)$. Then, each node has at least $2(r-1)$ neighbors. Thus, since $\mathcal{G}$ is an undirected graph, $|\mathcal{E}|\geq\frac{2n(r-1)}{2}=n(r-1)$.
Corollary 5.20 provides both the minimum degree required for each node and the total number of edges necessary for a graph to be $(r,r)$-robust. Similar to how we have used the notion of cliques as the building blocks of our analyses in Section 3, we will use Corollary 5.20 to analyze graphs with even $n$. By Corollary 5.20, we know for a graph with even $n$ to be $(\gamma,\gamma)$-MERG, it has to have at least $n(\gamma-1)=2\gamma(\gamma-1)$ edges. However, by leveraging the definition of $(r,s)$-robustness (Definition 8), we can derive a tighter lower bound:
Figure 3: This figure visualizes the contradiction shown in the proof for Lemma 5.22. The yellow solid lines represent edges, whereas the colored dotted lines represent the absences of edges. There are two disjoint non-empty subsets $\mathcal{S}_{1},\mathcal{S}_{2}\subset\mathcal{V}$, in which all nodes only have $\gamma-1$ neighbors outside of the subsets, necessitating at least $\lceil\frac{\gamma}{2}\rceil$ of the colored dotted lines to become actual edges to resolve the contradiction.
Lemma 5.22.
Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be a graph of $n$ nodes where $n$ is even, and let $\gamma=\frac{n}{2}$. Then, if $\mathcal{G}$ is $(\gamma,\gamma)$-robust,
$|\mathcal{E}|\geq 2\gamma(\gamma-1)+\Big{\lceil}{\frac{\gamma}{2}}\Big{\rceil}.$ (7)
Proof 5.23.
By Corollary 5.20, we know $\delta_{\min}(\mathcal{G})\geq 2(\gamma-1)$ and $|\mathcal{E}|\geq(2\gamma)(\gamma-1)$. In the rest of the proof, we will show that some nodes need more than $2(\gamma-1)$ neighbors for $\mathcal{G}$ to satisfy definition of $(r,s)$-robustness (Definition 8). Assume to the contrary that every node $i\in\mathcal{V}$ only has $2(\gamma-1)$ neighbors and $\mathcal{G}$ is $(\gamma,\gamma)$-robust. Because (i) there are $2\gamma$ nodes and (ii) each node has $2(\gamma-1)$ neighbors, we can form disjoint $\gamma$ pairs of nodes that do not have an edge together. Now, form a disjoint non-empty subsets $\mathcal{S}_{1},\mathcal{S}_{2}\subset\mathcal{V}$ such that each subset contains one node of each pair that do not have an edge. Since we have $\gamma$ pairs of nodes without an edge, $|\mathcal{S}_{1}|=|\mathcal{S}_{2}|=\gamma$. By the definition of $(r,s)$-robustness, we need at least $\gamma$ nodes that have $\gamma$ or more neighbors outside of their subsets. However, we have reached a contradiction, as all nodes in $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ have only $\gamma-1$ neighbors outside of their subsets. To resolve the contradiction, we need at least $\lceil\frac{\gamma}{2}\rceil$ edges, which are used to create edges between at least $\gamma$ nodes (as shown in Fig 3). Thus, combining everything, we need at least $2\gamma(\gamma-1)+\lceil\frac{\gamma}{2}\rceil$ edges.
Lemma 5.22 provides the necessary conditions on the edge counts of $(\gamma,\gamma)$-robust graphs with even numbers of nodes. We can also derive an interesting necessary graph structure by examining Lemma 5.22, which we provide below:
Corollary 5.24.
Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be a graph of $n$ nodes where $n$ is even, and let $\gamma=\frac{n}{2}$. Then, if $\mathcal{G}$ is $(\gamma,\gamma)$-robust, it must contain a $T(\gamma)$-clique, where
$T(\gamma)<\frac{2\gamma^{2}}{2\gamma-\big{\lceil}{\frac{\gamma}{2}}\big{\rceil}}+1.$ (8)
Proof 5.25.
We begin by recalling Turán’s theorem, which states that undirected graphs of $n$ nodes without a $k$-clique must have at most $(1-\frac{1}{k-1})\frac{n^{2}}{2}$ edges [43]. In other words, if a graph has more than $(1-\frac{1}{k-1})\frac{n^{2}}{2}$ edges, it must contain a $k$-clique.
Then, combining this with Lemma 5.22, $(\gamma,\gamma)$-robust graphs must contain a clique size of $T(\gamma)$, where
$2\gamma(\gamma-1)+\Big{\lceil}{\frac{\gamma}{2}}\Big{\rceil}$ $>\Big{(}1-\frac{1}{T(\gamma)-1}\Big{)}\frac{n^{2}}{2}$
$=2\gamma^{2}\Big{(}1-\frac{1}{T(\gamma)-1}\Big{)}.$
Rewriting the above, we get
$12\gamma-\Big{\lceil}{\frac{\gamma}{2}}\Big{\rceil}<\frac{2\gamma^{2}}{T(\gamma)-1}\to T(\gamma)<\frac{2\gamma^{2}}{2\gamma-\big{\lceil}{\frac{\gamma}{2}}\big{\rceil}}+$ 1 .
Corollary 5.24 characterizes the necessary maximum clique size that any $(\gamma,\gamma)$-robust graphs with even $n$ must contain. Roughly speaking, this implies the necessary maximum clique has to be at least size of $\lfloor{\frac{4\gamma}{3}}\rfloor+1$. Now, we will construct $(\gamma,\gamma)$-MERGs in the next section.
6 Construction of $(\gamma,\gamma)$-MERGs
Figure 4: Visualizations of $(5,5)$-MERGs with $9$ (left) and $10$ (right) nodes.
In the previous section, we have demonstrated the lower bounds on edge counts that any graph of $n$ nodes must satisfy in order to achieve $(\gamma,\gamma)$-robustness, where $\gamma=\lceil\frac{n}{2}\rceil$. Now, we present our second main result by constructing $(\gamma,\gamma)$-MERGs, or graphs that achieve $(\gamma,\gamma)$-robustness with a minimal set of edges:
Theorem 6.26.
Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be a graph of $n$ nodes, and let $\gamma=\lceil\frac{n}{2}\rceil$. Let every node have $2(\gamma-1)$ neighbors. If $n$ is even, find $\lceil\frac{\gamma}{2}\rceil$ pairs of nodes without an edge and connect them. Then, $\mathcal{G}$ is a $(\gamma,\gamma)$-MERG.
Proof 6.27.
Odd $n$: Because all $12\gamma-$ 1 nodes have $2(\gamma-1)$ neighbors, that means $\mathcal{G}$ forms a complete graph. Therefore by Lemma 5.18, $\mathcal{G}$ is a $(\gamma,\gamma)$-MERG when $n$ is odd.
Even $\mathbf{n}$: To prove that $\mathcal{G}$ is $(\gamma,\gamma)$-MERG for even values of $n$, we need to prove two things: (i) it is $(\gamma,\gamma)$-robust and (ii) its edge count $|\mathcal{E}|$ equals to the lower bounds specified in Lemma 5.22.
First, we prove its robustness. WLOG, we form disjoint non-empty subsets $\mathcal{S}_{1},\mathcal{S}_{2}\subset\mathcal{V}$ such that $|\mathcal{S}_{1}|\leq|\mathcal{S}_{2}|$, which implies $1\leq|\mathcal{S}_{1}|\leq\gamma$. By construction, there are at least $\gamma$ nodes in $\mathcal{G}$ that are connected to all other nodes (i.e., they has edges with $12\gamma-$ 1 nodes). There are two different cases:
1) If $|\mathcal{S}_{1}|\leq\gamma-1$, $|\mathcal{N}_{i}\setminus\mathcal{S}_{1}|\geq|\mathcal{N}_{i}|-\gamma-2\geq\gamma$ $\forall i\in\mathcal{S}_{1}$ since $|\mathcal{N}_{i}|\geq 2(\gamma-1)$ $\forall i\in\mathcal{V}$. Thus, every node in $\mathcal{S}_{1}$ has at least $\gamma$ neighbors outside of $\mathcal{S}_{1}$, which satisfies the first condition given in Definition 8.
2) If $|\mathcal{S}_{1}|=|\mathcal{S}_{2}|=\gamma$, note that we have at least $\gamma$ nodes that have $12\gamma-$ 1 neighbors. Thus, however we form $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$, there will be at least $\gamma=\frac{n}{2}$ nodes in $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ combined that have at least $2\gamma-1-(\gamma-1)=\gamma$ neighbors outside of $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$, respectively. This satisfies the third condition given in Definition 8. Since $\mathcal{G}$ satisfies Definition 8 with $r=s=\gamma$ in all cases, $\mathcal{G}$ is $(\gamma,\gamma)$-robust.
Now we show that $\mathcal{E}$ is a minimal set. We have a total degree of $4\gamma(\gamma-1)+K$, where $K=\gamma+1$ if $\gamma$ is odd and or $K=\gamma$ if $\gamma$ is even. Thus, the total number of edges is $|\mathcal{E}|=\frac{4\gamma(\gamma-1)+K}{2}=\frac{K}{2}+2\gamma(\gamma-1)$, which equals the lower bound given in Lemma 5.22.
Theorem 6.26 addresses Problem 2 by constructing $(\gamma,\gamma)$-MERGs, graphs that achieve maximum $(r,s)$-robustness with a minimal set of edges. Similar to $\gamma$-MERGs, these graphs also have as many edges as specified by the lower bounds in Lemma 5.18 and 5.22. This shows that the lower bounds given in the lemmas are tight. In fact, removing any edge from $(\gamma,\gamma)$-MERGs immediately reduces the graphs’ robustness levels likewise.
Examples of $(\gamma,\gamma)$-MERGs as the results of Theorem 6.26 are visualized in Figure 4. The graph on the left has $n=9$ nodes and forms a complete graph to achieve $(5,5)$-robustness, as discussed in Lemma 5.18. On the contrary, the graph on the right has $n=10$ nodes to achieve $(5,5)$-robustness. Note, it contains an $8$-clique as the maximum clique (i.e., $\{0,3,4,5,6,7,8,9\}$), which validates Corollary 5.24.
Furthermore, we also highlight that the construction mechanisms given in Theorem 6.26 are necessary and sufficient in a sense that they encompass all the possible variations of $(\gamma,\gamma)$-MERGs.
Proposition 1.
Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be a graph of $n$ nodes, and let $\gamma=\lceil\frac{n}{2}\rceil$. Then, $\mathcal{G}$ is $(\gamma,\gamma)$-robust if and only if $\mathcal{G}$ contains a graph $\mathcal{G}^{\prime}=(\mathcal{V},\mathcal{E}^{\prime})$ from Theorem 6.26 as a spanning subgraph (i.e., $\mathcal{E}^{\prime}\subseteq\mathcal{E}$).
Proof 6.28.
(necessary) Odd $n$: It is shown in Lemma 5.18 that complete graph is the only graph with $n$ nodes that is $(\gamma,s)$-robust where $s\geq\lfloor\frac{n}{2}\rfloor$. This indicates that $\mathcal{G}$ must be a complete graph, which means $\mathcal{E}^{\prime}=\mathcal{E}$, as $\mathcal{G}^{\prime}$ from Theorem 6.26 is also a complete graph.
Even $n$: Here we prove the claim by contradiction. Assume by contradiction that $\mathcal{G}$ is $(\gamma,\gamma)$-robust and $\mathcal{E}^{\prime}\not\subseteq\mathcal{E}$. Let $\mathcal{N}_{i}$ and $\mathcal{N}_{i}^{\prime}$ be the neighbor sets of agent $i\in\mathcal{V}$ in $\mathcal{G}$ and $\mathcal{G}^{\prime}$, respectively. Then, we define =$\mathcal{M}=\{i\in\mathcal{V}\mid|\mathcal{N}_{i}|=2\gamma-1\}$, $\mathcal{P}=\{i\in\mathcal{V}\mid|\mathcal{N}_{i}|=2\gamma-2\}$, =$\mathcal{M}^{\prime}=\{i\in\mathcal{V}\mid|\mathcal{N}_{i}^{\prime}|=2\gamma-1\}$, and $\mathcal{P}^{\prime}=\{i\in\mathcal{V}\mid|\mathcal{N}_{i}^{\prime}|=2\gamma-2\}$.
Note $\mathcal{G}^{\prime}$ has $|\mathcal{M}^{\prime}|=2\cdot\lceil\frac{\gamma}{2}\rceil$ nodes with degree of $12\gamma-$ 1 and $|\mathcal{P}^{\prime}|=2\gamma-|\mathcal{M}|$ nodes with degree of $22\gamma-$ 2. By Corollary 5.20, all nodes cannot have less than $22\gamma-$ 2 neighbors for $\mathcal{G}$ to be $(\gamma,\gamma)$-robust. Furthermore, they cannot have more than $12\gamma-$ 1 neighbors. What that means is that the only way for $\mathcal{E}^{\prime}\not\subseteq\mathcal{E}$ is to have $|\mathcal{M}|<|\mathcal{M}^{\prime}|$. Then, $|\mathcal{E}|<2\gamma(\gamma-1)+\lceil\frac{\gamma}{2}\rceil$. However, that is a contradiction, as for $\mathcal{G}$ to be $(\gamma,\gamma)$-robust, $|\mathcal{E}|\geq 2\gamma(\gamma-1)+\lceil\frac{\gamma}{2}\rceil$ by Lemma 5.22.
(sufficiency) We know that $\mathcal{G}^{\prime}$ is a $(\gamma,\gamma)$-MERG, i.e., $(\gamma,\gamma)$-robust graph with a minimal set of edges. Since $(r,s)$-robustness is monotonic with respect to number of edges [4, Lemma 3], if $\mathcal{E}^{\prime}\subseteq\mathcal{E}$, $\mathcal{G}$ must be $(\gamma,\gamma)$-robust.
Proposition 1 strengthens Theorem 6.26, as it demonstrates that the construction mechanism in the theorem captures all variations of possible $(\gamma,\gamma)$-MERGs. We can also conclude that if a graph does not contain $(\gamma,\gamma)$-MERGs from Theorem 6.26 as a spanning subgraph, it is not $(\gamma,\gamma)$-robust. In other words, we can use Proposition 1 to quickly verify whether a given graph $\mathcal{G}$ is $(\gamma,\gamma)$-robust by checking its spanning subgraphs. Although detecting such subgraphs is, in general, an NP-complete problem, many efficient subgraph matching algorithms exist (e.g., [44, 45]). This result can be then used to upper-bound the value of $F$, the maximum number of malicious agents a given graph can tolerate, by upper-bounding its $(r,s)$-robustness without using algorithms like [20, 23], which scale poorly with the number of nodes. This is particularly valuable, especially because computing $(r,s)$-robustness is generally more challenging than computing $r$-robustness.
Figure 5: Simulation results for the first set of simulations. We run consensus on $25$-MERGs with (a) $49$ and (b) $50$ nodes, and $(25,25)$-MERGs with (c) $49$ and (d) $50$ nodes, demonstrating resilient consensus through the W-MSR algorithm under $F=12$ and $F=24$ malicious agents, respectively. The normal agents’ states are solid colored lines, while malicious agents’ states are shown as red dotted lines.
Figure 6: Simulation results for the second set of simulations with $2$ Byzantine agents. We run consensus on $5$-MERGs with $n=9$ with and without an edge removal ((a) and (e)) and $n=10$ with and without an edge removal ((b) and (f)), and on $(5,5)$-MERGs with $n=9$ with and without an edge removal ((c) and (g)) and $n=10$ with and without an edge removal ((d) and (h)). The normal agents’ states are solid colored lines, while Byzantine agents’ states are shown as red dotted lines.
7 Simulations
In the previous sections, we have constructed $\gamma$- and $(\gamma,\gamma)$-MERGs, which provably achieve maximum $r$- and $(r,s)$-robustness with the fewest edges possible. In this section, we demonstrate the robustness of these graphs through two different sets of simulations (first in Sec. 7.1 and second in Sec. 7.2). We run each simulation from time step $t=0$ to $t=30$. In each simulation, the normals agents in a network run the W-MSR algorithm to perform consensus in the presence of $F$-total misbehaving agents, which transmit adversarial values to disrupt consensus. The simulation results are shown in Figure 5 and Figure 6, where the states of normal agents are plotted in solid colored lines, while states of misbehaving agents are plotted in red dotted lines.
7.1 Consensus Performance
In the first set of simulations, we demonstrate that in the presence of $F$ malicious agents, indexed by $i\in\{0,\dots,F-1\}$, normal agents $i\in\{F,\dots,n-1\}$ in $\gamma$- and $(\gamma,\gamma)$-MERGs are guaranteed to achieve resilient consensus. Each normal agent $i$’s initial state $x_{i}[0]\in\mathbb{R}$ is randomly generated on the interval $[-1000,1000]$. At time step $t\geq 0$, a malicious agent $i$ transmits an adversarial value: $1080\cos(t/5)$ if $i$ is even, and $1080\sin(t/5)$ if $i$ is odd.
We consider networks with $n=49$ and $n=50$ agents in all simulations. Maximum $r$-robustness any network with $49$ and $50$ agents can achieve is $r=\lceil\frac{n}{2}\rceil=25$-robustness. Thus, using Theorem 4.16, we construct $25$-MERGs with $49$ and $50$ nodes. Note these graphs achieve $25$-robustness with the fewest edges possible for a given number of nodes, and it can tolerate at most $F=12$ malicious agents. Thus, we run consensus simulations on these network graphs with $12$-total malicious agents, and the results are shown in Figure 5 (a) and (b). Similarly, the maximum $(r,s)$-robustness any network with $49$ and $50$ agents can achieve is $(25,25)$-robustness. Thus, we use Theorem 6.26 to construct $(25,25)$-MERGs with $49$ and $50$ agents. These graphs can tolerate up to $F=24$ malicious agents. We run consensus simulations on these network graphs with $24$-total malicious agents, and the results are shown in Figure 5 (c) and (d). As shown in Figure 5, both $\gamma$- and $(\gamma,\gamma)$-MERGs allow all normal agents to successfully achieve consensus within the convex hull of their initial values through the W-MSR, despite the presence of $F$ malicious agents.
7.2 Consensus with an Edge Removal
These $\gamma$- and $(\gamma,\gamma)$ graphs achieve their respective maximum robustness levels using the fewest edges possible for a given number of nodes. In other words, removing any edge reduces their robustness levels at least by one. To validate this minimalistic property, we compare the consensus performance of the original $\gamma$- and $(\gamma,\gamma)$-MERGs with that of modified versions in which a single edge has been removed. For each graph, comparison is conducted using identical initial states for the normal agents and in the presence of $F$ Byzantine agents.
7.2.1 Simulations with $\gamma$-MERGs
We construct $5$-MERGs with $n=9$ and $n=10$ nodes as shown in Figure 2. Note these graphs achieve $5$-robustness with the fewest edges possible for a given number of nodes, and they can tolerate at most $F=2$ byzantine agents. We set agents $i\in\{0,1\}$ as Byzantine, and agents $i\in\{2,\dots,n-1\}$ as normal. Each normal agent $i$ is initialized with a state $x_{i}[0]\in\mathbb{R}$ sampled uniformly at random from the interval $[15,100]$ for $i\in\{2,\dots,5\}$, from $[0,7]$ for $i\in\{6,\dots,n-2\}$, and from $[8,14]$ for $i=n-1$. For all $t\geq 0$, all Byzantine agents adversarially share a value of $100$ with agents $i=0,\dots,\lceil\frac{n}{2}\rceil$ and of 00 with other remaining agents.
Figure 6 (a)-(b) and (e)-(f) illustrate the consensus results with and without edge removal. We have removed the edge $(3,8)$ for $n=9$ and $(4,9)$ for $n=10$ from the $5$-MERGs shown in Figure 2. In the $5$-MERGs, normal agents successfully reach resilient consensus (shown in Figure 6 (e) and (f)). However, removing the edge $(3,8)$ from each $5$-MERG reduces its robustness to $r=4$, which is insufficient to guarantee resilient consensus in the presence of $F=2$ Byzantine agents. As a result, normal agents in the modified graph fail to reach resilient consensus (as shown by the red arrows in Figure 6 (a) and (b)).
7.2.2 Simulations with $(\gamma,\gamma)$-MERGs
We use Theorem 6.26 to construct $(5,5)$-MERGs with $n=9$ and $n=10$ agents, as shown in Figure 4. These graphs can tolerate up to $F=4$ Byzantine agents. We set agents $i\in\{0,\dots,3\}$ as Byzantine, and agents $i\in\{4,\dots,n-1\}$ as normal. Each normal agent $i$ is initialized with a state $x_{i}[0]\in\mathbb{R}$ sampled uniformly at random from the interval $[50,100]$ for $i\in\{4,\dots,n-2\}$ and from $[1,50]$ for $i=n-1$. Byzantine agents share adversarial values: agent $3$ transmits $100$ to its neighbors at all times, while the remaining Byzantine agents transmit 00 to their neighbors $\forall t\geq 0$.
Figure 6 (c)-(d) and (g)-(h) illustrate the consensus results with and without edge removal. We have removed the edge $(7,8)$ for $n=9$ and $(0,2)$ for $n=10$ from the $(5,5)$-MERGs visualized in Figure 4. In these $(5,5)$-MERGs, normal agents successfully reach resilient consensus (shown in Figure 6 (g) and (h)). However, removing the edge from each $(5,5)$-MERG reduces its robustness to at most $(5,4)$-robustness, which is insufficient to guarantee resilient consensus in the presence of $F=4$ Byzantine agents. As a result, normal agents fail to reach resilient consensus in the modified graph (as shown by the red arrows in Figure 6 (c) and (d)).
8 Conclusion
This paper investigates and presents the necessary conditions and topological structures of $r$- and $(r,s)$-robust graphs. These results are derived purely from the definitions of robustness, providing a general understanding of what these robustness levels require. We establish the tight lower bounds on the edge counts for graphs with any number of nodes to achieve maximum robustness. Then, we develop further from these results by constructing two classes of graphs of any number of nodes that achieve maximum $r$- and $(r,s)$-robustness with minimal sets of edges. For future work, we aim to extend our results to arbitrary robustness levels.
9 Appendix
Here, the proof for Lemma 3.7 is provided below, accompanied by Figure 7 for illustrative support.
Figure 7: These figures demonstrate the first two iterations of the first part in the proof of Lemma 3.7. Throughout figures (a) to (d), we maintain $|\mathcal{S}_{1}|=\gamma-1$ and $|\mathcal{S}_{2}|=\gamma+1$. Figures (a) and (b) illustrate the first iteration. Here, $\mathcal{S}_{2}$ contains $\gamma+1$ nodes including $k$-clique $\mathcal{C}_{k,1}=\{v_{1},\cdots,v_{k}\}$ (colored in green), while $\mathcal{S}_{1}$ contains the remaining $\gamma-1$ nodes. Since $\mathcal{S}_{1}$ has to be $\gamma$-reachable, there exists a node $u_{1}\in\mathcal{S}_{1}$ that has edges with at least $\gamma$ nodes in $\mathcal{S}_{2}$. Figure (a) shows the case scenario where the yellow node $u_{1}$ does not have an edge with $v_{1}$ (highlighted in yellow). Then, $u_{1}$ forms a $k$-clique $\mathcal{C}_{k,2}$ with $\mathcal{C}_{k,1}\setminus\{v_{1}\}$. Consequently, in any future iteration, if any node $u_{i}\in\mathcal{S}_{1}$, $i\neq 1$, connects to at least $\gamma$ nodes in $\mathcal{S}_{2}$ but not to $v_{1}$, it will form a $(k+1)$-clique with $\mathcal{C}_{k,2}$. Then, $u_{1}$ gets swapped with a node in $S_{2}\setminus\mathcal{C}_{k,1}$, as shown in Figure (b). Figures (c) and (d) present the second iteration. Figure (c) presents the case scenario where $u_{2}$ only has $\gamma$ edges and does not have an edge with $v_{2}$ (highlighted in red). As before, if any node $u_{i}\in\mathcal{S}_{2}$, $i\neq 2$, connects to $\gamma$ nodes while missing an edge to $v_{2}$ in future iteration, it will form a $(k+1)$-clique. Then, $u_{2}$ gets swapped with a node in $\mathcal{S}_{2}\setminus U^{\prime}$ where $U^{\prime}=\mathcal{C}_{k,1}\cup\mathcal{C}_{k,2}$, as shown on Figure (d). The process continues, maintaining $\gamma$-reachability of $\mathcal{S}_{1}$, until a $\big{(}\lfloor\frac{\gamma+3}{2}\rfloor\big{)}$-clique is formed.
Proof 9.29.
Since $\mathcal{G}$ is $\gamma$-robust, it holds that for any pair of disjoint, non-empty subsets $\mathcal{S}_{1}$, $\mathcal{S}_{2}\subset\mathcal{V}$, at least one of them is $\gamma$-reachable. The proof is divided into two parts. In the first part, we will inductively show our argument by constructing two subsets $\mathcal{S}_{1},\mathcal{S}_{2}\subset\mathcal{V}$ such that $\mathcal{G}$ is $\gamma$-robust and contains a $c$-clique where $2\leq c\leq\lfloor\frac{\gamma+3}{2}\rfloor$. Then, in the second part of the proof, we will show that $\mathcal{G}$ must contain a $\big{(}\lfloor\frac{\gamma+4}{2}\rfloor\big{)}$-clique through some reasoning.
First Part: For the induction step, we fix $|\mathcal{S}_{1}|=\gamma-1$ and $|\mathcal{S}_{2}|=\gamma+1$, enforcing $\mathcal{S}_{1}$ to be $\gamma$-reachable. Hence we have the base case of a $2$-clique. Now, we will iteratively swap nodes between $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ to argue some of the edges are necessary for $\mathcal{G}$ to satisfy Definition 6.
WLOG, let $\mathcal{S}_{2}$ contain $\gamma+1$ nodes including $k$-clique $\mathcal{C}_{k,1}=\{v_{1},\cdots,v_{k}\}$ (green in Figure 7), where $k\geq 2$, and let $\mathcal{S}_{1}$ contain the remaining $\gamma-1$ nodes. We also denote $U_{1}=U_{1}^{\prime}=\mathcal{C}_{k,1}$. Since $\mathcal{S}_{1}$ is $\gamma$-reachable, $\exists u_{1}\in\mathcal{S}_{1}$ (yellow in Figure 7) such that $|\mathcal{N}_{u_{1}}\cap\mathcal{S}_{2}|=|\mathcal{N}_{u_{1}}\setminus\mathcal{S}_{1}|\geq\gamma$. There are two cases:
  1. 1.
    $u_{1}$ has edges with all $\gamma+1$ nodes in $\mathcal{S}_{2}$, or $\gamma$ nodes but not with a node in $\mathcal{S}_{1}\setminus U_{1}$.
  2. 2.
    $u_{1}$ has edges with $\gamma$ nodes but not with $v_{1}\in U_{1}$ (highlighted in yellow in Figure 7).
If 1), $u_{1}$ forms a $(k+1)$-clique with $\mathcal{C}_{k,1}$. If 2), $u_{1}$ forms another $k$-clique $\mathcal{C}_{k,2}$ with $U_{1}^{\prime}\setminus\{v_{1}\}$. Note that $u_{1}$ has edges with all nodes in $\mathcal{S}_{2}\setminus\{v_{1}\}$. Now, if 1), we have a $(k+1)$-clique and done. If 2), let $U_{2}=\mathcal{C}_{k,1}\cap\mathcal{C}_{k,2}=\{v_{2},\cdots,v_{k}\}$ and $U_{2}^{\prime}=\mathcal{C}_{k,1}\cup\mathcal{C}_{k,2}=\{v_{1},\cdots,v_{k},u_{1}\}$. We also swap $u_{1}$ with any of the nodes in $\mathcal{S}_{2}\setminus U_{1}^{\prime}$ to get $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ shown on Figure 7 (b).
Then, since $\mathcal{S}_{1}$ has to be $\gamma$-reachable, $\exists u_{2}\in\mathcal{S}_{1}$ (red in Figure 7) such that $|\mathcal{N}_{u_{2}}\cap\mathcal{S}_{2}|\geq\gamma$. Again, there are two cases:
  1. 1.
    $u_{2}$ has edges with all $\gamma+1$ nodes in $\mathcal{S}_{2}$, or $\gamma$ nodes but not with a node in $\mathcal{S}_{2}\setminus U_{2}$
  2. 2.
    $u_{2}$ has edges with $\gamma$ nodes but not with $v_{2}\in U_{2}$ (highlighted in red in Figure 7).
If 1), $u_{2}$ forms a $(k+1)$-clique with $\mathcal{C}_{k,j}$ for some $j\in\{1,2\}$. If 2), $u_{2}$ forms another $k$-clique $\mathcal{C}_{k,3}$ with $U_{2}^{\prime}\setminus\{v_{1},v_{2}\}=\{v_{3},\cdots,v_{k},u_{1}\}$. Also note that $u_{2}$ has edges with all nodes in $\mathcal{S}_{2}\setminus\{v_{2}\}$ including $u_{1}$. Now, if 1), we have a $(k+1)$-clique and done. If 2), we update $U_{3}=\bigcap\limits_{j=1}^{3}\mathcal{C}_{k,j}=\{v_{3},\cdots,v_{k}\}$ and $U_{3}^{\prime}=\bigcup\limits_{j=1}^{3}\mathcal{C}_{k,j}=\{v_{1},\cdots,v_{k},u_{1},u_{2}\}$. We then swap $u_{2}$ with any node in $\mathcal{S}_{2}\setminus U_{2}^{\prime}$, continuing the process.
Likewise, unless we get case 1), we keep encountering the second case illustrated in the previous steps. To generalize the procedure, let $u_{n}\in\mathcal{S}_{1}$ such that $|\mathcal{N}_{u_{n}}\cap\mathcal{S}_{2}|\geq\gamma$, $n\in\{1,\cdots,k\}$. Then, $U_{n}=\bigcap\limits_{j=1}^{n}\mathcal{C}_{k,j}=\{v_{n},\cdots,v_{k}\}$ and $U_{n}^{\prime}=\bigcup\limits_{j=1}^{n}\mathcal{C}_{k,j}=\{v_{1},\cdots,v_{k},u_{1},\cdots,u_{n-1}$}. There are two cases:
  1. 1.
    $u_{n}$ has edges with all $\gamma+1$ nodes in $\mathcal{S}_{2}$, or $\gamma$ nodes but not with a node in $\mathcal{S}_{2}\setminus U_{n}$.
  2. 2.
    $u_{n}$ has edges with $\gamma$ nodes in $\mathcal{S}_{2}$ but not with $v_{n}\in U_{n}$.
If 1), $u_{n}$ forms a $(k+1)$-clique with $\mathcal{C}_{k,j}$ for any $j\in\{1,\cdots,n\}$. If 2), $u_{n}\in\mathcal{S}_{1}$ forms another $k$-clique $\mathcal{C}_{k,n+1}$ with $U_{n}^{\prime}\setminus\{v_{1},\cdots,v_{n}\}$. Now, if 1), we have a $(k+1)$-clique and done. If 2) we swap $u_{n}$ with any node in $\mathcal{S}_{2}\setminus U_{n}^{\prime}$ to continue the procedure.
If we continue encountering the case 2) for $k$ times, since $\mathcal{S}_{1}$ is $\gamma$-reachable, $\exists u_{k+1}\in\mathcal{S}_{1}$ such that $|\mathcal{N}_{u_{k+1}}\cap\mathcal{S}_{2}|\geq\gamma$. Then, there are two cases:
  1. 1.
    $u_{k+1}$ has edges with all $\gamma+1$ nodes in $\mathcal{S}_{2}$.
  2. 2.
    $u_{k+1}$ has edges with $\gamma$ nodes in $\mathcal{S}_{2}$ but not with a node in $\mathcal{S}_{2}\setminus U_{k+1}$ where $U_{k+1}=\bigcap\limits_{j=1}^{k+1}\mathcal{C}_{k,j}=\emptyset$.
In both cases, $u_{k+1}$ forms a $(k+1)$-clique with $\mathcal{C}_{k,j}$ for any $j\in\{1,\cdots,k+1\}$, showing a $(k+1)$-clique must exist. Also note that $U_{k+1}^{\prime}=\{v_{1},\cdots,v_{k},u_{1},\cdots,u_{k}\}\subseteq S_{2}$. Since (i) at worst $u_{1},\cdots,u_{k}$ as well as $v_{1},\cdots,v_{k}$ need to be in $\mathcal{S}_{2}$ for a $(k+1)$-clique to be formed and (ii) $|\mathcal{S}_{2}|=\gamma+1$, $k\leq\lfloor\frac{\gamma+1}{2}\rfloor$. That means $\mathcal{G}$ contains a $\big{(}\lfloor\frac{\gamma+3}{2}\rfloor\big{)}$-clique, which completes the first part of the proof.
Second Part: Continuing from the previous paragraph, at $k=\lfloor\frac{\gamma+1}{2}\rfloor$ and for even values of $\gamma$, at worst the case 2) mentioned above can be repeated $\frac{\gamma}{2}$ times. Then, $\mathcal{S}_{2}$ contains $\mathcal{K}=\{u_{1},\cdots,u_{\frac{\gamma}{2}}\}$, $\mathcal{C}_{k,1}=\{v_{1},\cdots,v_{\frac{\gamma}{2}}$}, and a node $p$ not in $\mathcal{K}$ and $\mathcal{C}_{k,1}$. Note both $\mathcal{K}$ and $\mathcal{C}_{k,1}$ disjointly form a $\big{(}\frac{\gamma}{2}\big{)}$-clique. We know that a node $u_{m}\in\mathcal{S}_{1}$, $m=\frac{\gamma+2}{2}$, must form a $m$-clique $\mathcal{Q}$ with either $\mathcal{K}$ or $\mathcal{C}_{k,1}$. We swap $u_{m}\in\mathcal{S}_{1}$ with $p\in\mathcal{S}_{2}$. Now let $\mathcal{P}=\mathcal{S}_{2}\setminus\mathcal{Q}$. Note that nodes in $\mathcal{P}$ form a $(\frac{\gamma}{2})$-clique such that $\mathcal{P}\cap\mathcal{Q}=\emptyset$. Since $\mathcal{S}_{1}$ is $\gamma$-reachable even after $u_{m}$ is swapped out, $\exists u_{m+1}\in\mathcal{S}_{1}$ such that $|\mathcal{N}_{u_{m+1}}\cap\mathcal{S}_{2}|\geq\gamma$. There are two cases:
  1. 1.
    $u_{m+1}$ has edges with all $\gamma+1$ nodes in $\mathcal{S}_{2}$, or $\gamma$ nodes but not with a node in $\mathcal{P}$.
  2. 2.
    $u_{m+1}$ has edges with $\gamma$ nodes in $\mathcal{S}_{2}$ but not with a node in $\mathcal{Q}$.
If 1), $\mathcal{Q}\cup\{u_{m+1}\}$ forms a $\big{(}\frac{\gamma+4}{2}\big{)}$-clique. If 2), $\mathcal{P}\cup\{u_{m+1}\}$ forms a $\big{(}\frac{\gamma+2}{2}\big{)}$-clique $\mathcal{P}^{\prime}$ such that $\mathcal{P}^{\prime}\cap\mathcal{Q}=\emptyset$. In this case, WLOG, let $\mathcal{P}^{\prime}\subset\mathcal{S}_{1}$ and $\mathcal{Q}\subset\mathcal{S}_{2}$ such that $|\mathcal{S}_{1}|=|\mathcal{S}_{2}|=\gamma$. Since either $\mathcal{S}_{1}$ or $\mathcal{S}_{2}$ is $\gamma$-reachable, all of the nodes in either $\mathcal{P}^{\prime}$ or $\mathcal{Q}$ have edges with one additional node in $\mathcal{S}_{2}$ or $\mathcal{S}_{1}$ respectively, forming a $\big{(}\frac{\gamma+4}{2}\big{)}$-clique. Thus, $\mathcal{G}$ must contain a $\big{(}\frac{\gamma+4}{2}\big{)}$-clique for even $\gamma$. Since $\mathcal{G}$ must have a clique size of $\frac{\gamma+3}{2}$ and $\frac{\gamma+4}{2}$ for odd and even $\gamma$ respectively, $\mathcal{G}$ must have a $\big{(}\lfloor{\frac{\gamma+4}{2}}\rfloor\big{)}$-clique.