### v2

# arXiv:2507.00415v2[eess.SY] 27 Feb 2026

Abstract
The notions of $r$-robustness and $(r,s)$-robustness of a network have been earlier introduced in the literature to achieve resilient consensus in the presence of misbehaving agents. However, while higher robustness levels enable networks to tolerate a higher number of misbehaving agents, they also require dense communication structures, which are not always desirable for systems with limited communication ranges, energy, and resources. Therefore, this paper studies the fundamental structures behind $r$-robustness and $(r,s)$- robustness properties in two ways. (a) We first establish tight necessary conditions on the number of edges that an undirected graph with an arbitrary number of nodes must have to achieve maximum $r$- and $(r,s)$-robustness. (b) We then use these conditions to construct two classes of undirected graphs, referred as to $\gamma$- and $(\gamma,\gamma)$-Minimal Edge Robust Graphs (MERGs), that provably achieve maximum robustness with minimal numbers of edges. We demonstrate the effectiveness of our method via comparison against existing robust graph structures and a set of simulations.
{IEEEkeywords}
Network control systems, distributed consensus and control, resilient consensus
1 Introduction
[…]
Distributed multi-agent systems have many practical uses, ranging from information gathering [1], target tracking [2], to collaborative decision making [3]. At the core of many distributed algorithms lies consensus, where multiple agents reach agreement on common state values. However, it is known that performance of consensus deteriorates significantly when one or more misbehaving agents share wrong, or even adversarial, information with the system. As a result, resilient consensus in the presence of misbehaving agents has been studied extensively [4, 5, 6, 7, 8, 9, 10, 11]. In [4, 6, 8], an algorithm called Weighted Mean-Subsequence-Reduced (W-MSR) algorithm was introduced to allow the non-misbehaving agents (often called normal) to reach consensus despite the presence of misbehaving agents. In contrast to reputation- and detection-based approaches [12, 13, 14] that require high computational overhead, the W-MSR algorithm instead ensures resilience through a low-complexity filtering strategy.
The notions of $r$- and $(r,s)$-robustness have been introduced in [4, 6] to characterize the resilience guarantees provided by the W-MSR algorithm. In particular, they quantify how many misbehaving agents the network tolerates while ensuring normal agents in a network to reach consensus within the convex hull of their initial values through the W-MSR algorithm [4]. $(2F+1)$-robustness provides a sufficient condition for normal agents to achieve consensus in the presence of $F$-local Byzantine agents. In contrast, $(F+1,F+1)$-robustness serves as a necessary and sufficient condition for tolerating $F$-total malicious agents [4]. Due to their distinct resilience guarantees, both $r$- and $(r,s)$-robustness have attracted significant attention and have been recently extended to ensure resilience in distributed estimation [15], optimization [16, 17], and learning [18, 19]. Their effectiveness has been demonstrated across a broad range of networked systems, including cooperative robotics and smart grids [20, 21, 22, 17].
Consequently, maximizing the robustness of a graph is highly desirable. However, achieving higher robustness by definition requires a higher number of edges, often referred to as edge counts. This results in dense and complex graph structures, which pose challenges for certain systems due to practical constraints such as limited communication range, bandwidth, and energy. This fundamental tradeoff between robustness and edge counts motivates us to investigate the following question: How can we systematically construct communication graphs that achieve maximum robustness with minimal edge counts?
Despite its clear practical importance, identifying the fundamental structures of robust graphs is particularly challenging due to the combinatorial nature of the problem. In fact, determining $r$-robustness is a coNP-complete problem [8]. Although algorithms for computing exact robustness exist [23, 24], they scale poorly with the number of nodes. This computational bottleneck makes brute-force search approaches impractical. Although there have been works to speed up computations with sampling-based [25], heuristics [26], or machine learning-based approaches [27], they either provide approximations without guarantees or require specific assumptions about graph topologies. In addition, other works under-bound a given graph’s robustness with other graph properties that are generally simpler and faster to compute [4, 22, 28, 20, 29]. However, these bounds are in general too loose for many graph structures.
To circumvent the combinatorial issue, some researches have focused on the construction and analysis of specific graph classes with predetermined $r$- and $(r,s)$-robustness properties. A complete graph with $n$ nodes is $\lceil\frac{n}{2}\rceil$-robust, indicating that $\lceil\frac{n}{2}\rceil$ is the maximum level of $r$-robustness any graph of $n$ nodes achieves [4]. Authors of [30] analyzed robustness of a class of circulant digraphs. A preferential-attachment method is investigated in [6, 4, 8] for scalable robust graph constructions. Furthermore, robustness levels of various graph structures, such as geometric and random graphs [31, 32, 8, 33, 34, 35] have been explored.
Still, understanding the fundamental structural properties of general $r$- and $(r,s)$-robust graphs remains an open problem. While the authors in [36] have established the lower bounds on edge counts and construction mechanisms for undirected $(2,2)$-robust graphs, these results do not generalize to other robustness levels. The authors in [37] study the construction of a class of undirected $r$-robust graphs with the minimum number of nodes. While this is equivalent to constructing graphs with maximum $r$-robustness for a given number of nodes, they do not provide any guarantees on the minimality of the edge set, and their construction method is valid only for specific values of $n$. The concept of minimal edge sets for $(r,s)$-robust graphs has been introduced in [38], though specific scalable construction mechanisms are not discussed.
Instead of investigating the fundamental structures of robust graphs, several studies have sought to relax the connectivity requirements associated with robustness. Some have explored time-varying network topologies to reduce connectivity requirements [39, 40, 41]. In [42], authors introduced trusted agents, who are always deemed fault-free, to achieve robustness with fewer edges. Furthermore, robustness under multi-hop communication schemes has been investigated in [43]. More recently, authors in [44] have proposed relaxing the connectivity requirements by employing detection-based isolation strategies. While these reduce the connectivity requirements associated with $r$- and $(r,s)$-robustness, they only do so by lowering the required robustness level and thus still depend on robust graph structures. This further underscores the need to investigate the fundamental structural properties of robustness.
Contributions
Our work directly investigates the fundamental structures of $r$- and $(r,s)$-robust graphs through two primary ways. First, we derive tight lower bounds on edge counts required for graphs with any number of nodes to achieve maximum robustness. To the best of our knowledge, these bounds are the first to explicitly depend on the level of robustness. Then, for an arbitrary number of nodes, we present two classes of graphs that provably achieve maximum robustness with the minimum number of edges. We refer to these graphs as $\gamma$- and $(\gamma,\gamma)$-Minimal Edge Robust Graphs (MERGs) [^1] In [45], we referred to graphs that achieve maximum $r$-robustness with minimal sets of edges as (2r1,r)(2r-1,r)- and $(2r,r)$-robust graphs. To avoid confusion with the standard definition of $(r,s)$-robustness, we adopt a new naming convention in this paper.. We build upon our previous work [45], in which we only studied $r$-robust graphs with minimal sets of edges, by also considering $(r,s)$-robust graphs. Specifically, we make the following contributions:
  1. 1.
    We establish tight lower bounds on the number of edges that an undirected graph with any number of nodes must have to achieve maximum $r$- or $(r,s)$-robustness.
  2. 2.
    Based on the lower bounds, we construct two classes of undirected graphs with an arbitrary number of nodes, referred to as $\gamma$- and $(\gamma,\gamma)$-MERGs that are proven to achieve maximum $r$- and $(r,s)$-robustness with minimal edge sets, respectively. Moreover, we prove that the construction mechanisms for $(\gamma,\gamma)$-MERGs are both necessary and sufficient, thereby capturing the entire space of graphs that achieve maximum $(r,s)$-robustness with the fewest edges.
  3. 3.
    We evaluate our methods through comparisons with other classes of robust graphs and a series of simulations.
The remainder of this paper is organized as follows. Section 2 presents preliminaries and the problem statement. Section 3 establishes necessary conditions on edge counts for graphs to achieve maximum $r$-robustness, whereas Section 4 provides systematic constructions of $\gamma$-MERGs. Section 5 determines the necessary conditions on edge counts for graphs to achieve maximum $(r,s)$-robustness, and Section 6 presents constructions of $(\gamma,\gamma)$-MERGs. Section 7 compares the proposed constructions with other existing graph constructions. Section 8 presents simulation results, and Section 9 concludes the paper and discusses directions for future research.
2 Background and Problem Statement
2.1 Notation
We denote a simple, undirected time-invariant graph as $\mathcal{G}=(\mathcal{V},\mathcal{E})$ where $\mathcal{V}$ and $\mathcal{E}$ are the vertex set and the undirected edge set of graph respectively. We denote the number of nodes $n$. An undirected edge $(i,j)\in\mathcal{E}$ indicates that information is exchanged between nodes $i$ and $j$. The neighbor set of agent $i\in\mathcal{V}$ is denoted as $\mathcal{N}_{i}=\{j\in\mathcal{V}\;|\;(i,j)\in\mathcal{E}\}$. The state of agent $i$ at time $t$ is denoted as $x_{i}[t]$. We denote the cardinality of a set $\mathcal{S}$ as $|\mathcal{S}|$. We denote the set of non-negative and positive integers as $\mathbb{Z}_{\geq 0}$ and $\mathbb{Z}_{>0}$. Lastly, we denote by $\lceil\cdot\rceil$ and $\lfloor\cdot\rfloor$ the ceiling and floor operators, respectively.
2.2 Weighted Mean Subsequence Reduced Algorithm
In this section, we review the relevant fundamental concepts of $r$- and $(r,s)$-robustness of a graph, as well as the W-MSR algorithm, which are defined in [4].
Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be a graph. Then, an agent $i\in\mathcal{V}$ shares its state $x_{i}[t]$ at time $t$ with all neighbors $j\in\mathcal{N}_{i}$, and each agent $i$ updates its state according to the nominal update rule:
$x_{i}[t+1]=w_{ii}[t]x_{i}[t]+\sum_{j\in\mathcal{N}_{i}}w_{ij}[t]x_{j}[t],$ (1)
where $w_{ij}[t]$ is the weight assigned to agent $j$’s value by agent $i$, and where the following conditions are assumed for $w_{ij}[t]$ for $\forall i\in\mathcal{V}$ and $t\in\mathbb{Z}_{\geq 0}$:
  • $w_{ij}[t]=0$ $\forall j\notin\mathcal{N}_{i}\cup\{i\}$,
  • $w_{ij}[t]\geq\alpha$, $0\leq\alpha<1$ $\forall j\in\mathcal{N}_{i}\cup\{i\}$,
  • $\sum_{j=1}^{n}w_{ij}[t]=1$
Through the protocol given by (1), agents are guaranteed to reach consensus on their states as long as the graph is connected [46, 47]. However, consensus may not be reached in the presence of misbehaving agents [4]. In our paper, we assume that the identities of misbehaving agents are unknown, and we focus on two types of misbehaving agents:
Definition 1 (Malicious agent).
An agent $i\in\mathcal{V}$ is called malicious if it does not follow the nominal update protocol (1) at some time step $t$.
Definition 2 (Byzantine agent).
An agent $i\in\mathcal{V}$ is called Byzantine if it does not follow the nominal update protocol (1), or if it does not send the same value to all of its neighbors at some time step $t$.
A malicious agent may deviate from the nominal update law but still broadcasts the same value to all of its neighbors. In contrast, a Byzantine agent may both deviate from the protocol and send different values to different neighbors. Hence, a malicious agent is as a special, weaker case of a Byzantine agent. Agents that are not misbehaving are called normal agents. There are numerous models to describe the number of misbehaving agents in a network [4]. We focus on two such models below:
Definition 3 ($\mathbf{F}$-total).
A set $\mathcal{S}\subset\mathcal{V}$ is $\mathbf{F}$-total if it contains at most $F$ nodes in the graph (i.e. $|\mathcal{S}|\leq F)$.
Definition 4 ($\mathbf{F}$-local).
A set $\mathcal{S}\subset\mathcal{V}$ is $\mathbf{F}$-local if all other nodes have at most $F$ nodes of $\mathcal{S}$ as their neighbors (i.e. $|\mathcal{N}_{i}\cap\mathcal{S}|\leq F$, $\forall i\in\mathcal{V}\setminus\mathcal{S}$).
In response to various threat and scope models, algorithms on resilient consensus [4, 39, 48, 5, 12, 13] have become very popular. In particular, the W-MSR (Weighted-Mean Subsequent Reduced) algorithm [4] with the parameter $F$ guarantees normal agents to achieve asymptotic consensus with $F$-total or $F$-local misbehaving agents under certain assumed topological properties of the communication graph.
Definition 5 ($\mathbf{r}$-reachable).
Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be a graph and $\mathcal{S}$ be a nonempty subset of $\mathcal{V}$. The subset $\mathcal{S}$ is $\mathbf{r}$-reachable if $\exists i\in\mathcal{S}$ such that $|\mathcal{N}_{i}\backslash\mathcal{S}|\geq r$, where $r\in\mathbb{Z}_{>0}$.
Definition 6 ($\mathbf{r}$-robust).
A graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ is $\mathbf{r}$-robust if $\forall\mathcal{S}_{1},\mathcal{S}_{2}\subset\mathcal{V}$ where $\mathcal{S}_{1}\cap S_{2}=\emptyset$ and $\mathcal{S}_{1},\mathcal{S}_{2}\neq\emptyset$, at least one of them is $r$-reachable.
Definition 7 ($\mathbf{(r,s)}$-reachable).
Given a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ and a nonempty subset $\mathcal{S}\subset\mathcal{V}$, $\mathcal{S}$ is $\mathbf{(r,s)}$-reachable if there are at least $s\in\mathbb{Z}_{\geq 0}$ nodes in $\mathcal{S}$ with at least $r\in\mathbb{Z}_{>0}$ neighbors outside of $\mathcal{S}$; i.e., given $\mathcal{X}^{r}_{\mathcal{S}}=\{i\in\mathcal{S}:|\mathcal{N}_{i}\setminus\mathcal{S}|\geq r\}$, then $|\mathcal{X}^{r}_{\mathcal{S}}|\geq s$.
Definition 8 ($\mathbf{(r,s)}$-robust).
A graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ is $\mathbf{(r,s)}$-robust if for every pair of nonempty, disjoint subsets $\mathcal{S}_{1},\mathcal{S}_{2}\subset\mathcal{V}$, at least one of the following holds (let $\mathcal{X}^{r}_{\mathcal{S}_{k}}=\{i\in\mathcal{S}_{k}:|\mathcal{N}_{i}\setminus\mathcal{S}_{k}|\geq r\}$ for $k\in\{1,2\}$):
  • $|\mathcal{X}^{r}_{\mathcal{S}_{1}}|=|\mathcal{S}_{1}|$;
  • $|\mathcal{X}^{r}_{\mathcal{S}_{2}}|=|\mathcal{S}_{2}|$;
  • $|\mathcal{X}^{r}_{\mathcal{S}_{1}}|$ + $|\mathcal{X}^{r}_{\mathcal{S}_{2}}|\geq s$.
A set $\mathcal{S}$ is $r$-reachable if it contains at least one node with at least $r$ neighbors outside the set, and $(r,s)$-reachable if it contains at least $s$ such nodes. These notions help us define $r$- and $(r,s)$-robustness. A graph is $r$-robust if, for every pair of disjoint subsets, at least one subset is $r$-reachable. In contrast, a graph is $(r,s)$-robust if, for every pair of disjoint subsets, at least $s$ nodes have at least $r$ neighbors outside their respective subsets. Hence, $(r,s)$-robustness generalizes $r$-robustness: any $(r,s)$-robust graph with $s\geq 1$ is also $r$-robust. Conversely, if a graph is $r$-robust, it is at least $(r,1)$-robust.
While $(2F+1)$-robustness provides a sufficient condition for consensus under $F$-local Byzantine agents, $(F+1,F+1)$-robustness is both necessary and sufficient condition to tolerate $F$-total malicious agents [4]. Note that these notions represent stronger forms of network resilience compared to more traditional graph-theoretic measures such as connectivity or minimum degree [4, 49]. In particular, $r$- and $(r,s)$-robust networks are also at least $r$-connected and has minimum degree of $r$ [4, 22]. Thus, these notions of robustness enable a network to achieve the same level of resilience (i.e., tolerance to the same number of misbehaving agents) using fewer edges (cf. [49, Fig. 3]).
2.3 Problem Statement
In this paper, we aim to construct graphs with a given number of nodes that achieve maximum $r$- and $(r,s)$- robustness using the minimal number of edges. Consider a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ where $|\mathcal{V}|=n$. Before formally defining our problems, we first define two classes of graphs below:
Definition 9 ($\mathbf{r}$-MERG).
A graph $\mathcal{G}$ is $\mathbf{r}$-Minimal Edge Robust Graph, or $\mathbf{r}$-MERG if $\mathcal{G}$ achieves $r$-robustness with the fewest edges possible.
Definition 10 ($\mathbf{(r,s)}$-MERG).
A graph $\mathcal{G}$ is $\mathbf{(r,s)}$-Minimal Edge Robust Graph, or $\mathbf{(r,s)}$-MERG if $\mathcal{G}$ achieves $(r,s)$-robustness with the fewest edges possible.
The $r$-MERGs and $(r,s)$-MERGs are general classes of graphs that achieve certain robustness levels with minimal sets of edges. Note that in [45] we have provided definitions of the same properties using different terms, which however may cause ambiguity and appear conflicting to the well-established notion of $(r,s)$-robust graphs. Hence in this paper we introduce the MERG definitions above to avoid any confusion. Since the maximum $r$-robustness $\mathcal{G}$ can achieve is $r=\lceil\frac{n}{2}\rceil$ [4], the problems addressed in this paper are:
Problem 1.
Given a desired number of nodes $n\in\mathbb{Z}_{>0}$, construct $\lceil\frac{n}{2}\rceil$-MERGs.
Problem 2.
Given a desired number of nodes $n\in\mathbb{Z}_{>0}$, construct $(\lceil\frac{n}{2}\rceil,\lceil\frac{n}{2}\rceil)$-MERGs.
Remark 1.
While the actual maximum $(r,s)$-robustness $\mathcal{G}$ achieves is $(\lceil\frac{n}{2}\rceil,n)$, this is only achieved by complete graphs [4, Lemma 4], which are rather trivial. Instead, we focus on the case where $r=s=\lceil\frac{n}{2}\rceil$, as $(\lceil\frac{n}{2}\rceil,\lceil\frac{n}{2}\rceil)$-robust network tolerates up to $\lceil\frac{n}{2}\rceil-1$ malicious agents, which is the maximum number of malicious agents the network with $n$ nodes tolerates through the W-MSR algorithm in practice.
For ease of notation, we define $\gamma=\lceil\frac{n}{2}\rceil$ for the remainder of the paper. Systematically constructing $\gamma$- and $(\gamma,\gamma)$-MERGs allows us to design networks that achieve the highest possible resilience for many distributed algorithms to misbehaving agents with the minimal number of communication links. Such constructions allow the system across diverse application domains to enjoy the highest robustness with reduced communication and energy overhead, which is critical for resource-constrained networks such as sensor networks, robotic swarms, and other distributed control systems. Moreover, the proposed constructions provide graph topologies that also work for extensions of the W-MSR algorithm which address practical network constraints, such as time-varying topologies and limited-rate communication [50, 39], as well as multi-dimensional states [51], without requiring additional edges.
Before proceeding with our analyses, we note that for both $\gamma$- and $(\gamma,\gamma)$-MERGs, we separate the analyses into two cases based on whether $n$ is even or odd. This separation is made for analytical convenience in our approach. For both cases, our analysis takes place in two steps: First, we establish the lower bounds on the edge counts for a graph with $n$ nodes to achieve maximum robustness ($\gamma$- and $(\gamma,\gamma)$-robustness). Second, we leverage these lower bounds to develop construction mechanisms that are proven to construct $\gamma$- and $(\gamma,\gamma)$-MERGs.
3 Bounds on Edge Counts for $\gamma$-MERGs
This section addresses the first step toward constructing $\gamma$-MERGs by examining the minimum number of edges required for a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ with $|\mathcal{V}|=n$ to achieve $\gamma=\lceil\frac{n}{2}\rceil$-robustness. This section consider two cases. In the first case, we find the lower bounds on edge counts for $\gamma$-robust graphs with odd $n$. Then, in the second case, we do the same for $\gamma$-robust graphs with even $n$. These are formally established in Lemma 2 and Lemma 5, respectively. These lemmas will be later used to construct $\gamma$-MERGs in Section 4.
Before analyzing the lower bounds on the edge counts for $\gamma$-robust graphs, we first introduce the concept of clique:
Definition 11 (clique [52]).
A clique $\mathcal{C}$ of a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ is a subset $\mathcal{C}\subseteq\mathcal{V}$ such that $\forall i,j\in\mathcal{C}$ with $i\neq j$, we have $(i,j)\in\mathcal{E}$.
In general, $k$-clique refers to a clique consisting of $k$ nodes. The clique with the largest cardinality in $\mathcal{G}$ is called the maximum clique. We will use cliques to discover necessary graph structures within $\gamma$-robust graphs, which will help us establish the lower bounds on edge counts.
3.1 Case 1: Graphs with Odd $n$
We now examine the lower bounds on the number of edges for the $\gamma$-robust graphs with odd $n$. To this end, we first study the smallest maximum clique that a $\gamma$-robust graph with odd $n$ must contain in order to satisfy the conditions of $r$-robustness (Definition 6).
Figure 1: These figures illustrate the first two iterations in the proof of Lemma 1. Figure (a) shows the initial configuration, where node $v_{1}\in\mathcal{V}$ has $\gamma$ neighbors (colored in yellow). Figure (b) shows the first iteration where node $v_{2}\in\mathcal{S}_{1}$ is connected to $\gamma-1$ purple nodes and to $v_{1}$ in $\mathcal{S}_{2}$. The green nodes $v_{1}$ and $v_{2}$ form a $2$-clique. Note we maintain $|\mathcal{S}_{1}|=\gamma-1$ and $|\mathcal{S}_{2}|=\gamma$ from figures (b) to (e). Figure (c) shows $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ after the completion of the first step, where $\gamma-1$ nodes in $\mathcal{S}_{1}$ (including $v_{2}$) have been swapped with $\gamma-1$ purple nodes in $\mathcal{S}_{2}\setminus\{v_{1}\}$. Then, the start of the second step is shown in Figure (d), where another node $v_{3}\in\mathcal{S}_{1}$ has edges with $\gamma-2$ red nodes and with $v_{1},v_{2}$ in $\mathcal{S}_{2}$. Here, the green nodes $v_{1},v_{2},$ and $v_{3}$ form a $3$-clique. Figure (e) shows $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ after $\gamma-2$ nodes in $\mathcal{S}_{1}$ (including $v_{3}$) have been swapped with $\gamma-2$ red nodes in $\mathcal{S}_{2}\setminus\{v_{1},v_{2}\}$. This process continues until $(\gamma+1)$-clique is formed.
Lemma 1.
Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be a graph of $n$ nodes where $n$ is odd, and let $\gamma=\lceil\frac{n}{2}\rceil$. Then, if $\mathcal{G}$ is $\gamma$-robust, it must contain a $(\gamma+1)$-clique.
Proof.
In this proof, we will show our argument by constructing two subsets $\mathcal{S}_{1},\mathcal{S}_{2}\subset\mathcal{V}$ such that $\mathcal{G}$ contains a $(\gamma+1)$-clique. Since $\mathcal{G}$ is $\gamma$-robust, it holds that for any pair of disjoint, non-empty subsets $\mathcal{S}_{1}$, $\mathcal{S}_{2}$ of $\mathcal{V}$, at least one of them is $\gamma$-reachable. Hence, by Definition 6 and [4, Lemma 5], for any node in $\mathcal{V}=\{v_{1},\dots,v_{n}\}$, $|\mathcal{N}_{v_{i}}|\geq\gamma$ must hold. WLOG, let $v_{1}\in\mathcal{V}$ be a node, as shown on Figure 1 (a). Then because $n=2\gamma-1$, we can construct $\mathcal{S}_{1}$ to be the set of $\gamma-1$ neighbors of $v_{1}$ and $\mathcal{S}_{2}$ to be the set of the remaining $\gamma$ nodes including $v_{1}$ such that $|\mathcal{S}_{1}|=\gamma-1$, $|\mathcal{S}_{2}|=\gamma$, and $\mathcal{S}_{1}\cap\mathcal{S}_{2}=\emptyset$. This enforces $\mathcal{S}_{1}$ to be $\gamma$-reachable.
Since $\mathcal{S}_{1}$ is $\gamma$-reachable, $\exists v_{2}\in\mathcal{S}_{1}$ such that $|\mathcal{N}_{v_{2}}\cap\mathcal{S}_{2}|=|\mathcal{N}_{v_{2}}\setminus\mathcal{S}_{1}|\geq\gamma$. Then in the first step, $v_{2}$ has an edge with $v_{1}\in\mathcal{S}_{2}$, forming a $2$-clique as shown on Figure 1 (b). We then swap $\gamma-1$ nodes in $\mathcal{S}_{1}$ including $v_{2}$ with $\gamma-1$ nodes in $\mathcal{S}_{2}\setminus\{v_{1}\}$ (colored in purple in Figure 1). Then in the second step, since $\mathcal{S}_{1}$ is $\gamma$-reachable even after $v_{2}$ is swapped out, $\exists v_{3}\in\mathcal{S}_{1}$ such that $|\mathcal{N}_{v_{3}}\cap\mathcal{S}_{2}|\geq\gamma$. By construction, since $v_{1}$ and $v_{2}$ are among the $\gamma$ nodes in $\mathcal{S}_{2}$, $v_{3}$ forms a $3$-clique with $v_{1}$ and $v_{2}$, as shown on Figure 1 (d). We then swap $\gamma-2$ nodes in $\mathcal{S}_{1}$ including $v_{3}$ with $\gamma-2$ nodes in $\mathcal{S}_{2}\setminus\{v_{1},v_{2}\}$ (colored in red in Figure 1). Again, in the third step, since $\mathcal{S}_{1}$ is $\gamma$-reachable even after $v_{3}$ is swapped out, $\exists v_{4}\in\mathcal{S}_{1}$ that has edges with all $\gamma$ nodes in $\mathcal{S}_{2}$. Since $v_{1},v_{2},v_{3}\in\mathcal{S}_{2}$ are three of the $\gamma$ nodes to have edges with $v_{4}$, the nodes $v_{1},v_{2},v_{3},v_{4}$ form a $4$-clique. We then swap $\gamma-3$ nodes in $\mathcal{S}_{1}$ including $v_{4}$ with $\gamma-3$ nodes in $\mathcal{S}_{2}\setminus\{v_{1},v_{2},v_{3}\}$, and thus the construction continues.
Likewise, in the $k^{\text{th}}$ step, $\exists v_{k+1}\in\mathcal{S}_{1}$ such that $|\mathcal{N}_{v_{k+1}}\cap\mathcal{S}_{2}|\geq\gamma$ and thus forms a $(k+1)$-clique with $v_{1},v_{2},\cdots,v_{k-1},v_{k}\in\mathcal{S}_{2}$. We then swap $\gamma-k$ nodes in $\mathcal{S}_{1}$ including $v_{k+1}$ with $\gamma-k$ nodes in $\mathcal{S}_{2}\setminus\{v_{1},\cdots,v_{k}\}$. At the end of $\gamma-1^{\text{th}}$ step, $\mathcal{S}_{2}$ contains $v_{1},\cdots,v_{\gamma}$ that form a $\gamma$-clique. Finally, since $\mathcal{S}_{1}$ is $\gamma$-reachable, it contains a node $v_{\gamma+1}$ that has edges with all nodes $v_{1},\cdots,v_{\gamma}\in\mathcal{S}_{2}$. Then, $\{v_{\gamma+1}\}\cup\mathcal{S}_{2}$ forms a ($\gamma+1)$-clique. ∎
Lemma 1 states that for a graph with odd $n$ to be $\gamma$-robust, it must contain a clique size of $\gamma+1$ or more nodes. Stemming from this reasoning, we finalize the lower bound on edge counts for the case of odd $n$:
Lemma 2.
Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be a graph of $n$ nodes where $n$ is odd, and let $\gamma=\lceil\frac{n}{2}\rceil$. Then, if $\mathcal{G}$ is $\gamma$-robust,
$|\mathcal{E}|\geq\frac{3\gamma(\gamma-1)}{2}.$ (2)
Proof.
Since $\mathcal{G}$ is $\gamma$-robust, it holds that for any pair of disjoint, non-empty subsets $\mathcal{S}_{1}$, $\mathcal{S}_{2}\subset\mathcal{V}$, at least one of them is $\gamma$-reachable (by Definition 6). From Lemma 1, we know that $\mathcal{V}$ must contain a $(\gamma+1)$-clique $\mathcal{C}$. This means that $\mathcal{G}$ must have at least as many edges as to construct a $(\gamma+1)$-clique, i.e., $|\mathcal{E}|\geq\frac{\gamma(\gamma+1)}{2}$.
Let $\mathcal{C}^{\prime}=\mathcal{V}\setminus\mathcal{C}$. This then implies $|\mathcal{C}|=\gamma+1$ and $|\mathcal{C}^{\prime}|=\gamma-2$. WLOG, let $\mathcal{S}_{1}=\mathcal{C}^{\prime}$ and $\mathcal{S}_{2}=\mathcal{C}$. Then, $|\mathcal{S}_{1}|=\gamma-2$ and $|\mathcal{S}_{2}|=\gamma+1$. This implies $\mathcal{S}_{1}$ has to be $\gamma$-reachable. In other words, there must exist $v_{1}\in\mathcal{S}_{1}$ such that $|\mathcal{N}_{v_{1}}\cap\mathcal{S}_{2}|=|\mathcal{N}_{v_{1}}\setminus\mathcal{S}_{1}|\geq\gamma$. Therefore, for such $v_{1}$ to exist, we need $\gamma$ edges that connect $v_{1}$ with $\gamma$ nodes in $\mathcal{S}_{2}$: $|\mathcal{E}|\geq\frac{\gamma(\gamma+1)}{2}+\gamma$. Note that $v_{1}\in\mathcal{S}_{1}$ is the node that makes $\mathcal{S}_{1}$ $\gamma$-reachable. Therefore, moving $v_{1}$ from $\mathcal{S}_{1}$ to $\mathcal{S}_{2}$ requires the new $\mathcal{S}_{1}$, which is now without $v_{1}$, to contain another node $v_{2}$ that has at least $\gamma$ neighbors in $\mathcal{S}_{2}$. This again requires at least $\gamma$ additional edges (i.e. $|\mathcal{E}|\geq\frac{\gamma(\gamma+1)}{2}+2\gamma$). We then move $v_{2}$ from $\mathcal{S}_{1}$ to $\mathcal{S}_{2}$, which necessitates $\mathcal{S}_{1}$ to have another node $v_{3}$ to have $\gamma$ neighbors in $\mathcal{S}_{2}$.
We continue this process of (i) connecting a node $v_{i}\in\mathcal{S}_{1}$ with any $\gamma$ nodes in $\mathcal{S}_{2}$ and (ii) moving $v_{i}$ from $\mathcal{S}_{1}$ to $\mathcal{S}_{2}$ in the $i^{\text{th}}$ step until $\mathcal{S}_{1}$ becomes empty. Then, it is possible to repeat these two steps $\gamma-2$ times, as we initially had $|\mathcal{S}_{1}|=\gamma-2$. In other words, for each $\gamma-2$ pair of $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$, we require $\gamma$ additional edges. Since we started from the setting where $\mathcal{S}_{2}=\mathcal{C}$ and forced $|\mathcal{E}|$ to increase at least a total of $\gamma(\gamma-2)$, $|\mathcal{E}|\geq\frac{\gamma(\gamma+1)+2\gamma(\gamma-2)}{2}\Rightarrow|\mathcal{E}|\geq\frac{3\gamma(\gamma-1)}{2}$. ∎
Using Lemma 2, we conclude that any $\gamma$-robust graphs with odd $n$, i.e., graphs with an odd number of nodes that achieve maximum $r$-robustness, must have $\frac{3\gamma(\gamma-1)}{2}$ or more edges. In fact, Theorem 1 later shows that the bound in the lemma is tight and cannot be further improved. Furthermore, using Lemma 2, we derive a useful corollary:
Corollary 1.
Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be an $r$-robust graph, where $r\in\mathbb{Z}_{>0}$. Then,
$|\mathcal{E}|\geq\frac{3r(r-1)}{2}.$ (3)
Proof.
In general, the more nodes a graph has, the more edges it must have to maintain a given robustness. From [37], it is shown that 2r12r-1 is the smallest number of nodes a graph must have to be $r$-robust. In other words, $r$-robustness is the maximum robustness a graph with 2r12r-1 nodes achieves. Thus, Lemma 2 is restated as: any $r$-robust graph with 2r12r-1 nodes must have at least $\frac{3r(r-1)}{2}$ edges. Consequently, any graph $\mathcal{G}$ with 2r12r-1 or more nodes must have at least $\frac{3r(r-1)}{2}$ edges to be $r$-robust. ∎
Remark 2.
Corollary 1 reveals that the bound in Lemma 2 is in fact a necessary condition for any $r$-robust graphs of any number of nodes $n$, ensuring that graphs with edges less than $\frac{3r(r-1)}{2}$ cannot be $r$-robust. One useful implication of this result is that it provides a quick way to upper-bound a graph’s robustness level based solely on its number of edges. Directly computing $r$-robustness of a graph is generally difficult [8, 24], and while various approximation methods exist [25, 26, 27], they often lack formal guarantees or assume some structures on the given graphs. In fact, the upper bound on robustness derived from this result can be used to accelerate existing methods by reducing their search space.
3.2 Case 2: Graphs with Even $n$
Now we continue our analysis on the edge counts for the graphs with even $n$. Since $n$ is even, $\gamma=\frac{n}{2}$. This implies that, for the same value of $\gamma$, the value of $n$ is one greater when $n$ is even compared to when it is odd. However, the overall approach is similar to the previous case: we will first determine the minimum size of the maximum clique that $\mathcal{G}$ must contain to be $\gamma$-robust.
Lemma 3.
Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be a graph with $n$ nodes where $n$ is even, and let $\gamma=\frac{n}{2}$. Then, if $\mathcal{G}$ is $\gamma$-robust, it must contain a $\big(\lfloor\frac{\gamma+4}{2}\rfloor\big)$-clique.
Due to length, the proof for Lemma 3 is given in the Appendix (Section 10). We emphasize that the bounds from Lemmas 1 and 3 are only analytical lower bounds we are able to show, and may not reflect the actual necessary maximum clique size. In fact, unlike the analysis for odd $n$, studying $\gamma$-robust graphs with even $n$ involves topological structures beyond the maximum clique we observe in Lemma 3. The following lemma presents the additional necessary structures.
Lemma 4.
Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be a graph of $n$ nodes where $n$ is even, and let $\gamma=\frac{n}{2}$. Then, if $\mathcal{G}$ is $\gamma$-robust, $\mathcal{G}$ contains an induced subgraph $\mathcal{G}_{S}=(\mathcal{V}_{S},\mathcal{E}_{S})$ such that $|\mathcal{V}_{S}|=\gamma+1$ and $|\mathcal{E}_{S}|\geq\lfloor{\frac{\gamma^{2}+2}{2}}\rfloor$.
Proof.
Since $\mathcal{G}$ is $\gamma$-robust, for any pair of disjoint, non-empty subsets $\mathcal{S}_{1}$, $\mathcal{S}_{2}\subset\mathcal{V}$, at least one of them is $\gamma$-reachable. From Lemma 3, $\mathcal{G}$ contains a $\big(\lfloor\frac{\gamma+4}{2}\rfloor\big)$-clique $\mathcal{C}$. WLOG, let $\mathcal{S}_{2}$ contain $\gamma+1$ nodes including $\mathcal{C}$, and let $\mathcal{S}_{1}$ contain the remaining $\gamma-1$ nodes. Initially, we set $\mathcal{V}_{S}=\mathcal{C}$. Then, $|\mathcal{E}_{S}|=\big(\frac{1}{2}\big)\big(\lfloor\frac{\gamma+4}{2}\rfloor\big)\big(\lfloor\frac{\gamma+2}{2}\rfloor\big)$. Note there are total $n=2\gamma$ nodes. Since $|\mathcal{S}_{1}|=\gamma-1$ and $|\mathcal{S}_{2}|=\gamma+1$, $\mathcal{S}_{1}$ must be $\gamma$-reachable.
By Lemma 3, we have a $2$-clique and $3$-clique when $\gamma=1$ and $\gamma=2$ respectively, satisfying the statement in the lemma for $\gamma=1,2$.
Now, we consider $\gamma\geq 3$. With this setup, since $\mathcal{S}_{1}$ is $\gamma$-reachable, $\exists v_{1}\in\mathcal{S}_{1}$ such that $|\mathcal{N}_{v_{1}}\cap\mathcal{S}_{2}|=|\mathcal{N}_{v_{1}}\setminus\mathcal{S}_{1}|\geq\gamma$. We swap $v_{1}\in\mathcal{S}_{1}$ with any node in $\mathcal{S}_{2}\setminus\mathcal{V}_{S}$. Note that $|\mathcal{V}_{S}|=\lfloor\frac{\gamma+4}{2}\rfloor$ at this point. Now, we add $v_{1}$ as a vertex of $\mathcal{G}_{S}$ (i.e. $\mathcal{V}_{S}$ now contains $v_{1}$). Then, $|\mathcal{E}_{S}|$ increases at least by $\lfloor\frac{\gamma+2}{2}\rfloor$, since $v_{1}$ may have no edge with one of the nodes in $\mathcal{V}_{S}$. Since $\mathcal{S}_{1}$ is $\gamma$-reachable even without $v_{1}$, $\exists v_{2}\in\mathcal{S}_{1}$ such that $|\mathcal{N}_{v_{2}}\cap\mathcal{S}_{2}|\geq\gamma$. Note that $|\mathcal{V}_{S}|=\lfloor\frac{\gamma+6}{2}\rfloor$ at this point. We swap $v_{2}\in\mathcal{S}_{1}$ with a node in $\mathcal{S}_{2}\setminus\mathcal{V}_{S}$ and add $v_{2}$ as a new vertex of $\mathcal{G}_{S}$ (i.e. $\mathcal{V}_{S}$ now contains $v_{2}$). This increases $|\mathcal{E}_{S}|$ at least by $\lfloor\frac{\gamma+4}{2}\rfloor$, as $v_{2}$ may have no edge with one of the nodes in $\mathcal{V}_{S}$. Then, since $\mathcal{S}_{1}$ is still $\gamma$-reachable, $\exists v_{3}\in\mathcal{S}_{1}$ such that $|\mathcal{N}_{v_{3}}\cap\mathcal{S}_{2}|\geq\gamma$. Note that $|\mathcal{V}_{S}|=\lfloor\frac{\gamma+8}{2}\rfloor$. Again, we swap $v_{3}\in\mathcal{S}_{1}$ with a node in $\mathcal{S}_{2}\setminus\mathcal{V}_{S}$ and add $v_{3}$ as a new vertex of $\mathcal{G}_{S}$. Then, $|\mathcal{E}_{S}|$ increases at least by $\lfloor\frac{\gamma+6}{2}\rfloor$.
This process continues until when $v_{p}\in\mathcal{S}_{1}$, $p=\gamma+1-\lfloor\frac{\gamma+4}{2}\rfloor$, that makes $\mathcal{S}_{1}$ $\gamma$-reachable and gets swapped with a node in $\mathcal{S}_{2}\setminus\mathcal{V}_{S}$ where $\mathcal{V}_{S}=\mathcal{C}\cup\{v_{1},v_{2},\cdots,v_{p-1}\}$. Here, adding $v_{p}$ as a new vertex of $\mathcal{G}_{S}$ increases $|\mathcal{E}_{S}|$ at least by $\gamma-1$. At this point, $\mathcal{S}_{2}=\mathcal{V}_{S}$ and $|\mathcal{V}_{S}|=\gamma+1$. Then, $|\mathcal{E}_{S}|\geq\big(\frac{1}{2}\big)\big(\lfloor\frac{\gamma+4}{2}\rfloor\big)\big(\lfloor\frac{\gamma+2}{2}\rfloor\big)+(\lfloor\frac{\gamma+2}{2}\rfloor)+(\lfloor\frac{\gamma+4}{2}\rfloor)+(\lfloor\frac{\gamma+6}{2}\rfloor)+\cdots+(\gamma-2)+(\gamma-1)$.
We remove the floor functions: if $\gamma$ is even, $|\mathcal{E}_{S}|\geq\big(\frac{1}{2}\big)\big(\frac{\gamma+4}{2}\big)\big(\frac{\gamma+2}{2}\big)+\big(\frac{\gamma+2}{2}\big)+\big(\frac{\gamma+4}{2}\big)+\cdots+(\gamma-2)+(\gamma-1)=\frac{\gamma^{2}+2}{2}$. If $\gamma$ is odd, $|\mathcal{E}_{S}|\geq\big(\frac{1}{2}\big)\big(\frac{\gamma+3}{2}\big)\big(\frac{\gamma+1}{2}\big)+\big(\frac{\gamma+1}{2}\big)+\big(\frac{\gamma+3}{2}\big)+\cdots+(\gamma-2)+(\gamma-1)=\frac{\gamma^{2}+1}{2}$. Thus, we have $|\mathcal{V}_{S}|=\gamma+1$ and $|\mathcal{E}_{S}|\geq\lfloor{\frac{\gamma^{2}+2}{2}}\rfloor$. ∎
Lemma 4 effectively provides a necessary induced subgraph structure that $\mathcal{G}$ with even $n$ must have to be $\gamma$-robust. Now, using this lemma, we present the following:
Lemma 5.
Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be a graph of $n$ nodes where $n$ is even, and let $\gamma=\frac{n}{2}$. Then, if $\mathcal{G}$ is $\gamma$-robust,
$|\mathcal{E}|\geq\left\lfloor{\frac{\gamma(3\gamma-2)+2}{2}}\right\rfloor.$ (4)
Proof.
The approach of the proof is identical to that of Lemma 2, except for the fact that in this proof, we are starting with an induced subgraph $\mathcal{G}_{S}=(\mathcal{V}_{S},\mathcal{E}_{S})$ such that $|\mathcal{V}_{S}|=\gamma+1$ and $|\mathcal{E}_{S}|\geq\lfloor{\frac{r^{2}+2}{2}}\rfloor$ (which comes from Lemma 4).
Let $\mathcal{S}_{2}=\mathcal{V}_{S}$ and $\mathcal{S}_{1}=\mathcal{V}\setminus\mathcal{V}_{S}$, which means $|\mathcal{S}_{1}|=\gamma-1$ and $|\mathcal{S}_{2}|=\gamma+1$. This enforces $\mathcal{S}_{1}$ to be $\gamma$-reachable. Then, $\exists v_{1}\in\mathcal{S}_{1}$ such that $|\mathcal{N}_{v_{1}}\cap\mathcal{S}_{2}|=|\mathcal{N}_{v_{1}}\setminus\mathcal{S}_{1}|\geq\gamma$. This requires at least $\gamma$ edges (i.e. $|\mathcal{E}|\geq|\mathcal{E}_{S}|+\gamma$). We move $v_{1}$ from $\mathcal{S}_{1}$ to $\mathcal{S}_{2}$. Then since $\mathcal{S}_{1}$ is $\gamma$-reachable even without $v_{1}$, $\exists v_{2}\in\mathcal{S}_{1}$ such that $|\mathcal{N}_{v_{2}}\cap\mathcal{S}_{2}|\geq\gamma$. This requires at least $\gamma$ new edges (i.e. $|\mathcal{E}|\geq|\mathcal{E}_{S}|+2\gamma$).
We continue this process of (i) drawing edges between $v_{i}\in\mathcal{S}_{1}$ and $\gamma$ nodes in $\mathcal{S}_{2}$ and (ii) moving $v_{i}\in\mathcal{S}_{1}$ into $\mathcal{S}_{2}$ until $\mathcal{S}_{1}$ becomes empty. Then, we can do these steps $\gamma-1$ times, as we initially have $|\mathcal{S}_{1}|=\gamma-1$. In other words, for each $\gamma-1$ pair of $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$, we require $\gamma$ additional edges. Since we have started from the setting where $\mathcal{S}_{2}=\mathcal{V}_{S}$, we have $|\mathcal{E}|\geq|\mathcal{E}_{S}|+\gamma(\gamma-1)$. Therefore, if $\gamma$ is even, $|\mathcal{E}|\geq\frac{\gamma^{2}+2}{2}+\gamma(\gamma-1)=\frac{\gamma(3\gamma-2)+2}{2}$. If $\gamma$ is odd, and $|\mathcal{E}|\geq\frac{\gamma^{2}+1}{2}+\gamma(\gamma-1)=\frac{\gamma(3\gamma-2)+1}{2}$. ∎
[…]
4 Construction of $\gamma$-MERGs
In the previous section, we have established the lower bounds on edge counts that graphs of $n$ nodes must satisfy in order to achieve $\gamma$-robustness (where $\gamma=\lceil\frac{n}{2}\rceil$ is the maximum robustness). Proceeding with the lower bounds, we construct $\gamma$-MERGs, or graphs that provably achieve $\gamma$-robustness with a minimal set of edges. This is formally given in Theorem 1, as our first main result:
Theorem 1.
Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be a graph with $n$ nodes, and let $\gamma=\lceil\frac{n}{2}\rceil$. Then, $\mathcal{G}$ is a $\gamma$-MERG under the following construction mechanisms:
  • If $n$ is odd, let $\mathcal{X}\subset\mathcal{V}$ be a set of $\gamma+1$ nodes that form a $(\gamma+1)$-clique. Connect each of the remaining $\gamma-1$ nodes to any $\gamma$ nodes in $\mathcal{X}$.
  • If $n$ is even, let $\mathcal{X}\subset\mathcal{V}$ be a set of $\gamma$ nodes, each of which is adjacent to all other nodes in $\mathcal{V}$. Select $\lceil\frac{\gamma-2}{2}\rceil$ disjoint pairs of nodes in $\mathcal{X}$ and remove the edge between each pair.
Proof.
To prove that $\mathcal{G}$ is $\gamma$-MERG, we need to prove two things for both odd and even values of $n$: (i) it is $\gamma$-robust and (ii) its edge count $|\mathcal{E}|$ equals to the lower bounds specified in Lemma 2 or Lemma 5, depending on the value of $n$.
Odd $\mathbf{n}$: We first prove that it is $\gamma$-robust. Let $\mathcal{X}^{\prime}=\mathcal{V}\setminus\mathcal{X}$. Then, $|\mathcal{X}|=\gamma+1$ and $|\mathcal{X}^{\prime}|=\gamma-2$. WLOG, let $\mathcal{S}_{1},\mathcal{S}_{2}\in\mathcal{V}$ be nonempty subsets such that $\mathcal{S}_{1}\cap\mathcal{S}_{2}=\emptyset$ and $|\mathcal{S}_{1}|\leq|\mathcal{S}_{2}|$. There are two cases.
1) If $\mathcal{X}\cap\mathcal{S}_{1}=\emptyset$, let a node $i\in\mathcal{X}^{\prime}\cap\mathcal{S}_{1}$. Since $\mathcal{X}\cap\mathcal{S}_{1}=\emptyset$ and $\mathcal{N}_{i}\subset\mathcal{X}\subseteq\mathcal{S}_{2}$ by construction, $|\mathcal{N}_{i}\setminus\mathcal{S}_{1}|\geq\gamma$, making $\mathcal{S}_{1}$ $\gamma$-reachable.
2) If $\mathcal{X}\cap\mathcal{S}_{1}\neq\emptyset$, denote $|\mathcal{X}\cap\mathcal{S}_{1}|=p\leq\gamma-1$. Then, we have $1\leq|\mathcal{X}^{\prime}\cap\mathcal{S}_{1}|\leq\gamma-1-p$, $|\mathcal{X}\setminus\mathcal{S}_{1}|=\gamma+1-p$, and $p-1\leq|\mathcal{X}^{\prime}\setminus\mathcal{S}_{1}|\leq\gamma-2$. For $i\in\mathcal{X}\cap\mathcal{S}_{1}$, $|(\mathcal{N}_{i}\setminus\mathcal{S}_{1})\cap\mathcal{X}|=|\mathcal{X}\setminus\{i\}|-(p-1)=\gamma-(p-1)$, as $\mathcal{X}$ forms a $(\gamma+1)$-clique. For $p=1$, $\exists i\in\mathcal{X}\cap\mathcal{S}_{1}$ such that $|\mathcal{N}_{i}\setminus\mathcal{S}_{1}|\geq|\mathcal{X}\setminus\{i\}|=\gamma$. For $p\geq 2$, note that by construction, every node in $\mathcal{X}^{\prime}$ has edges with any $\gamma$ nodes in $\mathcal{X}$, where $|\mathcal{X}|=\gamma+1$. Speaking differently, every node in $\mathcal{X}^{\prime}$ does not have an edge with only one node in $\mathcal{X}$. Thus, all nodes in $\mathcal{X}^{\prime}\setminus\mathcal{S}_{1}$ have edges with at least p1p-1 nodes in $\mathcal{X}\cap\mathcal{S}_{1}$, because it cannot have an edge with only one node in $\mathcal{X}\cap\mathcal{S}_{1}$. Then, as $p-1\leq|\mathcal{X}^{\prime}\setminus\mathcal{S}_{1}|\leq\gamma-2$, there are at least a total of (p1)(p1)$(p-1)$ edges between the nodes in $\mathcal{X}^{\prime}\setminus\mathcal{S}_{1}$ and $\mathcal{X}\cap\mathcal{S}_{1}$. Since $|\mathcal{X}\cap\mathcal{S}_{1}|=p$ and by Pigeonhole Principle, $\exists i\in\mathcal{X}\cap\mathcal{S}_{1}$ such that $|(\mathcal{N}_{i}\setminus\mathcal{S}_{1})\cap\mathcal{X}^{\prime}|\geq\big\lceil{\frac{(p-1)(p-1)}{p}}\big\rceil=\big\lceil{\frac{p^{2}-2p+1}{p}}\big\rceil=p-1$. Then, $\exists i\in\mathcal{X}\cap\mathcal{S}_{1}$ such that $|\mathcal{N}_{i}\setminus\mathcal{S}_{1}|=|(\mathcal{N}_{i}\setminus\mathcal{S}_{1})\cap\mathcal{X}|+|(\mathcal{N}_{i}\setminus\mathcal{S}_{1})\cap\mathcal{X}^{\prime}|\geq\gamma-(p-1)+p-1=\gamma$. Thus, whether $\mathcal{S}_{1}$ contains a node in $\mathcal{X}$ or not, $\mathcal{S}_{1}$ is always $\gamma$-reachable.
Now we show that $\mathcal{E}$ is a minimal set. $\mathcal{G}$ has $\frac{(\gamma+1)(\gamma)}{2}$ edges to form a $(\gamma+1)$-clique. Additionally, we introduce $\gamma(\gamma-2)$ edges for connecting each of the $\gamma-2$ nodes in $\mathcal{X}^{\prime}$ with $\gamma$ nodes in $\mathcal{X}$. Summing them up, we have $\frac{3\gamma(\gamma-1)}{2}$ which equals the lower bound given in Lemma 2.
Even $\mathbf{n}$: We first prove that it is $\gamma$-robust. Let $\mathcal{X}^{\prime}=\mathcal{V}\setminus\mathcal{X}$. Then, $|\mathcal{X}|=\gamma$ and $|\mathcal{X}^{\prime}|=\gamma$. WLOG, let $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ be nonempty subsets of $\mathcal{V}$ such that $\mathcal{S}_{1}\cap\mathcal{S}_{2}=\emptyset$ and $|\mathcal{S}_{1}|\leq|\mathcal{S}_{2}|$. That means $1\leq|\mathcal{S}_{1}|\leq\gamma$. There are two cases.
1) If $\mathcal{X}\cap\mathcal{S}_{1}=\emptyset$, let a node $i\in\mathcal{X}^{\prime}\cap\mathcal{S}_{1}$. Since $\mathcal{X}\cap\mathcal{S}_{1}=\emptyset$ and $\mathcal{N}_{i}=\mathcal{X}$, $|\mathcal{N}_{i}\setminus\mathcal{S}_{1}|\geq\gamma$, and thus $\mathcal{S}_{1}$ is $\gamma$-reachable.
2) If $\mathcal{X}\cap\mathcal{S}_{1}\neq\emptyset$, let a node $i\in\mathcal{X}\cap\mathcal{S}_{1}$. Note that after removing an edge among $\lceil\frac{\gamma-2}{2}\rceil$ pairs of nodes in $\mathcal{X}$, there exists at least one (two in case of even $\gamma$) node(s) that is connected to all other nodes in $\mathcal{G}$. In other words, $\exists v_{*}\in\mathcal{V}$ s.t. $|N_{v_{*}}|=2\gamma-1$. Now, when $1\leq|\mathcal{S}_{1}|\leq\gamma-1$, $|\mathcal{N}_{i}\setminus\mathcal{S}_{1}|\geq\gamma$, making $\mathcal{S}_{1}$ $\gamma$-reachable. In case $|\mathcal{S}_{1}|=|\mathcal{S}_{2}|=\gamma$, we know there exists $v_{*}$ in either $\mathcal{S}_{1}$ or $\mathcal{S}_{2}$ such that $|\mathcal{N}_{v_{*}}\setminus\mathcal{S}_{i}|\geq\gamma$, making $\mathcal{S}_{i}$ $\gamma$-reachable. Thus, whether $\mathcal{S}_{1}$ contains a node in $\mathcal{X}$ or not, either $\mathcal{S}_{1}$ or $\mathcal{S}_{2}$ is always $\gamma$-reachable.
Now, we examine its edge set. The $\gamma$ nodes in $\mathcal{X}$ connecting to every other node adds up to a total of $\frac{\gamma(\gamma-1)+2\gamma(\gamma)}{2}=\frac{3\gamma^{2}-\gamma}{2}$ edges. Then, if we subtract $\lceil\frac{\gamma-2}{2}\rceil$ from it, we get $\frac{\gamma(3\gamma-2)+2}{2}$ for even $\gamma$ and $\frac{\gamma(3\gamma-\gamma)+1}{2}$ for odd $\gamma$, which equals the lower bound given in Lemma 5. ∎
Figure 2: Visualizations of $5$-MERGs with $9$ (left) and $10$ (right) nodes.
Theorem 1 provides a systematic procedure for constructing $\gamma$-MERGs, i.e., graphs that achieve maximum $r$-robustness with a minimal set of edges, thereby solving Problem 1. Examples of $\gamma$-MERGs with $n=9$ (left) and $n=10$ (right) are visualized in Figure 2. The theorem also suggests that the bounds given in Lemma 2 and 5 are tight. In fact, removing any edge from a $\gamma$-MERG immediately drops its robustness by at least one, as demonstrated later in Section 8.2.
Remark 3.
We observe that the construction of $F$-elemental graphs in [37, Prop. 1] bears similarities to our approach in Theorem 1 for odd $n$. Both methods aim to achieve maximum $r$-robustness and include a $\gamma$-clique; the key difference lies in how the remaining non-clique nodes are connected. We also observe that the total number of edges required to achieve maximum robustness for odd $n$ is the same for both approaches (see Table 1), suggesting the existence of a more general construction framework for odd $n$. While generalizing these constructions is left for future work, we note that our method additionally handles even $n$ and provides a formal proof of edge-count minimality, which is not addressed in [37].
5 Bounds on Edge Counts for $(\gamma,\gamma)$-MERGs
In Sections 3 and 4, we have examined and constructed $\gamma$-MERGs to address Problem 1. In the following two sections, we focus on $(\gamma,\gamma)$-MERGs, graphs with a given number of nodes that achieve maximum $(r,s)$-robustness with a minimal set of edges, thereby addressing Problem 2.
Similar to what we have done in the previous sections, we consider two cases: odd $n$ and even $n$. However, note that $(\gamma,\gamma)$-MERGs for odd $n$ have already been characterized in [4, Lemma 4], which is given below for completeness:
Lemma 6.
[4] Let $n>1$ be odd and $\gamma=\lceil\frac{n}{2}\rceil$. Then, complete graphs are the only graphs with $n$ nodes that are $(\gamma,s)$-robust with $s\geq\lfloor\frac{n}{2}\rfloor$.
Lemma 6 essentially confirms that when $n$ is odd, the graph must be a complete graph to be a $(\gamma,\gamma)$-MERG. This also implies that when $n$ is odd, for a graph to be $(\gamma,\gamma)$-robust, it must have at least $\frac{(2\gamma-1)(2\gamma-2)}{2}=(\gamma-1)(2\gamma-1)$ edges.
5.1 Graphs with Even $n$
We now introduce the lower bounds on edge counts for $(\gamma,\gamma)$-robust graphs when $n$ is even. This is formally established in Lemma 8. This lemma will be later used in Section 6 to construct $(\gamma,\gamma)$-MERGs for any $n$. However, before that, we introduce another useful lemma from [4]:
Lemma 7.
Given an $(r,s)$-robust graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, with $0\leq r\leq\lceil\frac{n}{2}\rceil$ and $1\leq s\leq n$, the minimum degree of $\mathcal{G}$, $\delta_{\min}(\mathcal{G})$, is at least
$\delta_{\min}(\mathcal{G})\geq\begin{cases}2r-2&\text{if }s\geq r\\ r+s-1&\text{otherwise}\end{cases}.$ (5)
Because we consider $r=s=\gamma$ in our paper, what we really care in (5) is the inequality for the first case: $\delta_{\min}(\mathcal{G})\geq 2r-2$. Therefore, we have the following corollary:
Corollary 2.
Given an $(r,r)$-robust graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ of $n$ nodes, the minimum degree of $\mathcal{G}$, $\delta_{\min}(\mathcal{G})\geq 2(r-1)$. Furthermore, $|\mathcal{E}|\geq n(r-1)$.
Proof.
As the direct application of Equation 5, we conclude $\delta_{\min}(\mathcal{G})\geq 2(r-1)$. Then, each node has at least 2(r1)2(r-1) neighbors. Thus, since $\mathcal{G}$ is an undirected graph, $|\mathcal{E}|\geq\frac{2n(r-1)}{2}=n(r-1)$. ∎
Corollary 2 provides both the minimum degree required for each node and the total number of edges necessary for a graph to be $(r,r)$-robust. Similar to how we have used the notion of cliques as the building blocks of our analyses in Section 3, we will use Corollary 2 to analyze graphs with even $n$. By Corollary 2, we know for a graph with even $n$ to be $(\gamma,\gamma)$-MERG, it must have at least $n(\gamma-1)=2\gamma(\gamma-1)$ edges. However, by leveraging the definition of $(r,s)$-robustness (Definition 8), we derive a tighter lower bound:
Figure 3: This figure visualizes the contradiction shown in the proof for Lemma 8. The yellow solid lines represent edges, whereas the colored dotted lines represent the absences of edges. There are two disjoint non-empty subsets $\mathcal{S}_{1},\mathcal{S}_{2}\subset\mathcal{V}$, in which all nodes only have $\gamma-1$ neighbors outside of the subsets, necessitating at least $\lceil\frac{\gamma}{2}\rceil$ of the colored dotted lines to become actual edges to resolve the contradiction.
Lemma 8.
Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be a graph of $n$ nodes where $n$ is even, and let $\gamma=\frac{n}{2}$. Then, if $\mathcal{G}$ is $(\gamma,\gamma)$-robust,
$|\mathcal{E}|\geq 2\gamma(\gamma-1)+\Big\lceil{\frac{\gamma}{2}}\Big\rceil.$ (6)
Proof.
By Corollary 2, we know $\delta_{\min}(\mathcal{G})\geq 2(\gamma-1)$ and $|\mathcal{E}|\geq(2\gamma)(\gamma-1)$. In the rest of the proof, we will show that some nodes need more than $2(\gamma-1)$ neighbors for $\mathcal{G}$ to satisfy definition of $(r,s)$-robustness (Definition 8). Assume to the contrary that every node $i\in\mathcal{V}$ only has $2(\gamma-1)$ neighbors and $\mathcal{G}$ is $(\gamma,\gamma)$-robust. Because (i) there are $2\gamma$ nodes and (ii) each node has $2(\gamma-1)$ neighbors, it is possible to form disjoint $\gamma$ pairs of nodes that do not have an edge together. Now, form a disjoint non-empty subsets $\mathcal{S}_{1},\mathcal{S}_{2}\subset\mathcal{V}$ such that each subset contains one node of each pair that do not have an edge. Since we have $\gamma$ pairs of nodes without an edge, $|\mathcal{S}_{1}|=|\mathcal{S}_{2}|=\gamma$. By the definition of $(r,s)$-robustness, we need at least $\gamma$ nodes that have $\gamma$ or more neighbors outside of their subsets. However, we have reached a contradiction, as all nodes in $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ have only $\gamma-1$ neighbors outside of their subsets. To resolve the contradiction, we need at least $\lceil\frac{\gamma}{2}\rceil$ edges, which are used to create edges between at least $\gamma$ nodes (as shown in Fig 3). Thus, combining everything, we need at least $2\gamma(\gamma-1)+\lceil\frac{\gamma}{2}\rceil$ edges. ∎
Lemma 8 provides the necessary conditions on the edge counts of $(\gamma,\gamma)$-robust graphs with even numbers of nodes. Combining Lemmas 6 and 8, we will construct $(\gamma,\gamma)$-MERGs in the next section.
6 Construction of $(\gamma,\gamma)$-MERGs
Figure 4: Visualizations of $(5,5)$-MERGs with $9$ (left) and $10$ (right) nodes.
In the previous section, we presented the lower bounds on edge counts that any graph with $n$ nodes must satisfy in order to achieve $(\gamma,\gamma)$-robustness, where $\gamma=\lceil\frac{n}{2}\rceil$. Now, we present our second main result by constructing $(\gamma,\gamma)$-MERGs, or graphs that achieve $(\gamma,\gamma)$-robustness with a minimal set of edges:
Theorem 2.
Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be a graph of $n$ nodes, and let $\gamma=\lceil\frac{n}{2}\rceil$. Let every node have $2(\gamma-1)$ neighbors. If $n$ is even, find $\lceil\frac{\gamma}{2}\rceil$ pairs of nodes without an edge and connect them. Then, $\mathcal{G}$ is a $(\gamma,\gamma)$-MERG.
Proof.
Odd $n$: Because all $2\gamma-1$ nodes have $2(\gamma-1)$ neighbors, that means $\mathcal{G}$ forms a complete graph. Therefore by Lemma 6, $\mathcal{G}$ is a $(\gamma,\gamma)$-MERG when $n$ is odd.
Even $\mathbf{n}$: To prove that $\mathcal{G}$ is $(\gamma,\gamma)$-MERG for even values of $n$, we need to prove two things: (i) it is $(\gamma,\gamma)$-robust and (ii) its edge count $|\mathcal{E}|$ equals to the lower bounds specified in Lemma 8.
First, we prove its robustness. WLOG, we form disjoint non-empty subsets $\mathcal{S}_{1},\mathcal{S}_{2}\subset\mathcal{V}$ such that $|\mathcal{S}_{1}|\leq|\mathcal{S}_{2}|$, which implies $1\leq|\mathcal{S}_{1}|\leq\gamma$. By construction, there are at least $\gamma$ nodes in $\mathcal{G}$ that are connected to all other nodes (i.e., they has edges with $2\gamma-1$ nodes). There are two different cases:
1) If $|\mathcal{S}_{1}|\leq\gamma-1$, $|\mathcal{N}_{i}\setminus\mathcal{S}_{1}|\geq|\mathcal{N}_{i}|-\gamma-2\geq\gamma$ $\forall i\in\mathcal{S}_{1}$ since $|\mathcal{N}_{i}|\geq 2(\gamma-1)$ $\forall i\in\mathcal{V}$. Thus, every node in $\mathcal{S}_{1}$ has at least $\gamma$ neighbors outside of $\mathcal{S}_{1}$, which satisfies the first condition given in Definition 8.
2) If $|\mathcal{S}_{1}|=|\mathcal{S}_{2}|=\gamma$, note that we have at least $\gamma$ nodes that have $2\gamma-1$ neighbors. Thus, however we form $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$, there will be at least $\gamma=\frac{n}{2}$ nodes in $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ combined that have at least $2\gamma-1-(\gamma-1)=\gamma$ neighbors outside of $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$, respectively. This satisfies the third condition given in Definition 8. Since $\mathcal{G}$ satisfies Definition 8 with $r=s=\gamma$ in all cases, $\mathcal{G}$ is $(\gamma,\gamma)$-robust.
Now we show that $\mathcal{E}$ is a minimal set. We have a total degree of $4\gamma(\gamma-1)+K$, where $K=\gamma+1$ if $\gamma$ is odd and or $K=\gamma$ if $\gamma$ is even. Thus, the total number of edges is $|\mathcal{E}|=\frac{4\gamma(\gamma-1)+K}{2}=\frac{K}{2}+2\gamma(\gamma-1)$, which equals the lower bound given in Lemma 8. ∎
Theorem 2 addresses Problem 2 by constructing $(\gamma,\gamma)$-MERGs, graphs that achieve maximum $(r,s)$-robustness with a minimal set of edges. Similar to $\gamma$-MERGs, these graphs also have as many edges as specified by the lower bounds in Lemmas 6 and 8. This shows that the lower bounds given in the lemmas are tight. In fact, removing any edge from $(\gamma,\gamma)$-MERGs immediately reduces the graphs’ robustness and thus resilience to misbehaving agents.
Examples of $(\gamma,\gamma)$-MERGs as the results of Theorem 2 are visualized in Figure 4. The graph on the left has $n=9$ nodes and forms a complete graph to achieve $(5,5)$-robustness, as discussed in Lemma 6. On the contrary, the graph on the right has $n=10$ nodes and is $(5,5)$-robust.
Furthermore, we highlight that the construction mechanisms given in Theorem 2 are necessary and sufficient in a sense that they encompass all the possible variations of $(\gamma,\gamma)$-MERGs.
Proposition 1.
Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be a graph of $n$ nodes, and let $\gamma=\lceil\frac{n}{2}\rceil$. Then, $\mathcal{G}$ is $(\gamma,\gamma)$-robust if and only if $\mathcal{G}$ contains a graph $\mathcal{G}^{\prime}=(\mathcal{V},\mathcal{E}^{\prime})$ from Theorem 2 as a spanning subgraph (i.e., $\mathcal{E}^{\prime}\subseteq\mathcal{E}$).
Proof.
(necessary) Odd $n$: It is shown in Lemma 6 that complete graph is the only graph with $n$ nodes that is $(\gamma,s)$-robust where $s\geq\lfloor\frac{n}{2}\rfloor$. This indicates that $\mathcal{G}$ must be a complete graph, which means $\mathcal{E}^{\prime}=\mathcal{E}$, as $\mathcal{G}^{\prime}$ from Theorem 2 is also a complete graph.
Even $n$: Here we prove the claim by contradiction. Assume by contradiction that $\mathcal{G}$ is $(\gamma,\gamma)$-robust and $\mathcal{E}^{\prime}\not\subseteq\mathcal{E}$. Let $\mathcal{N}_{i}$ and $\mathcal{N}_{i}^{\prime}$ be the neighbor sets of agent $i\in\mathcal{V}$ in $\mathcal{G}$ and $\mathcal{G}^{\prime}$, respectively. Then, we define $\mathcal{M}=\{i\in\mathcal{V}\mid|\mathcal{N}_{i}|=2\gamma-1\}$, $\mathcal{P}=\{i\in\mathcal{V}\mid|\mathcal{N}_{i}|=2\gamma-2\}$, $\mathcal{M}^{\prime}=\{i\in\mathcal{V}\mid|\mathcal{N}_{i}^{\prime}|=2\gamma-1\}$, and $\mathcal{P}^{\prime}=\{i\in\mathcal{V}\mid|\mathcal{N}_{i}^{\prime}|=2\gamma-2\}$.
Note $\mathcal{G}^{\prime}$ has $|\mathcal{M}^{\prime}|=2\cdot\lceil\frac{\gamma}{2}\rceil$ nodes with degree of $2\gamma-1$ and $|\mathcal{P}^{\prime}|=2\gamma-|\mathcal{M}|$ nodes with degree of $2\gamma-2$. By Corollary 2, all nodes cannot have less than $2\gamma-2$ neighbors for $\mathcal{G}$ to be $(\gamma,\gamma)$-robust. Furthermore, they cannot have more than $2\gamma-1$ neighbors. What that means is that the only way for $\mathcal{E}^{\prime}\not\subseteq\mathcal{E}$ is to have $|\mathcal{M}|<|\mathcal{M}^{\prime}|$. Then, $|\mathcal{E}|<2\gamma(\gamma-1)+\lceil\frac{\gamma}{2}\rceil$. However, that is a contradiction, as for $\mathcal{G}$ to be $(\gamma,\gamma)$-robust, $|\mathcal{E}|\geq 2\gamma(\gamma-1)+\lceil\frac{\gamma}{2}\rceil$ by Lemma 8.
(sufficiency) We know that $\mathcal{G}^{\prime}$ is a $(\gamma,\gamma)$-MERG, i.e., $(\gamma,\gamma)$-robust graph with a minimal set of edges. Since $(r,s)$-robustness is monotonic with respect to number of edges [4, Lemma 3], if $\mathcal{E}^{\prime}\subseteq\mathcal{E}$, $\mathcal{G}$ must be $(\gamma,\gamma)$-robust. ∎
It follows from 1 that every $(\gamma,\gamma)$-MERG - that is, every $(\gamma,\gamma)$-robust graph with a minimal edge set - can be constructed using the procedure in Theorem 2. Consequently, it is possible to use 1 to quickly verify whether a given graph $\mathcal{G}$ is $(\gamma,\gamma)$-robust by checking whether it contains $(\gamma,\gamma)$-MERGs from Theorem 2 as a spanning subgraph. While finding such subgraphs is in general NP-complete, many efficient algorithms exist (e.g., [53, 54]).
Figure 5: Simulation results for the first set of simulations. We run consensus on $25$-MERGs with (a) $49$ and (b) $50$ nodes, and $(25,25)$-MERGs with (c) $49$ and (d) $50$ nodes, demonstrating resilient consensus through the W-MSR algorithm under $F=12$ and $F=24$ malicious agents, respectively. The normal agents’ states are solid colored lines, while malicious agents’ states are shown as red dotted lines.
Figure 6: Simulation results for the second set of simulations. Consensus with $F=2$ Byzantine agents is performed on 55-MERGs with $n=9$ (panels (a) and (e)) and $n=10$ (panels (b) and (f)), both without and with an edge removal, respectively. Similarly, consensus with $F=4$ malicious agents is executed on $(5,5)$-MERGs with $n=9$ (panels (c) and (g)) and $n=10$ (panels (d) and (h)), also without and with an edge removal, respectively. The states of normal agents are shown as solid lines, while those of misbehaving agents are omitted for visual clarity.
7 Comparison with Other Graphs
Table 1: Number of edges required to achieve maximum robustness for different graph construction methods as the number of nodes $n$ increases.
$n$ Graphs Edge Count Edge Count
($\gamma$-robustness) ($(\gamma,\gamma)$-robustness)
5 $k$-Circulant [49] 10 10
$F$-Elemental [37] 9 -
Ours 9 10
30 $k$-Circulant [49] 435 435
$F$-Elemental [37] - -
Ours 323 428
105 $k$-Circulant [49] 5460 5460
$F$-Elemental [37] 4134 -
Ours 4134 5460
300 $k$-Circulant [49] 44850 44850
$F$-Elemental [37] - -
Ours 33601 44775
In this section, we compare $\gamma$- and $(\gamma,\gamma)$-MERGs from Theorems 1 and 2 with other existing robust graphs in terms of edge counts. The goal is to highlight that our construction methods require less edges than other existing methods. We focus on two classes of graphs for comparison: undirected $k$-circulant graphs [49, 8] and $F$-elemental graphs [37]. Table 1 summarizes the number of edges required to achieve maximum robustness for each class of graphs as the number of nodes $n$ increases.
Undirected $k$-circulant graph is $\lceil k/2\rceil$-robust and at least $(\lfloor(k+2)/2\rfloor,\lfloor(k+2)/2\rfloor)$-robust for even $k$ and $(\lfloor(k+1)/2\rfloor,\lfloor(k+1)/2\rfloor)$-robust for odd $k$ [49, 55]. Consequently, achieving the maximum robustness level requires the graph to be fully connected, leading to a high number of edges, as shown in Table 1. On the other hand, $F$-Elemental graphs constructed with [37, Prop. 1] (discussed in Remark 3) tend to use as few edges as ours. However, these graphs are defined only $r$-robust graphs with odd $n$ and provide no analysis of edge minimality. In contrast, as supported by our analyses in Sections 3 and 5, our graphs constructed according to Theorems 1 and 2 consistently use the fewest edges for any $n$ and for both $r$- and $(r,s)$-robustness.
8 Simulations
In the previous sections, we have established systematic ways to construct $\gamma$- and $(\gamma,\gamma)$-MERGs, which provably achieve maximum $r$- and $(r,s)$-robustness with the fewest edges possible. We now validate the correctness of our results through two sets of simulations (presented in Section 8.1 and Section 8.2). While $\gamma$- and $(\gamma,\gamma)$-MERGs support a variety of resilient algorithms, including estimation, optimization, and learning [15, 16, 18] as well as different resilient consensus algorithms [43, 39, 56, 57], here we focus specifically on resilient consensus via the W-MSR algorithm.
For all simulations, normal agents in the network execute the W-MSR algorithm to achieve consensus despite the presence of misbehaving agents, which transmit adversarial values to disrupt consensus. Through the W-MSR algorithm, normal agents’ states are guaranteed to converge to a common value within the convex hull of their initial conditions, irrespective of the behaviors of the misbehaving agents [4]. The results are shown in Figures 5 and 6.
8.1 Consensus Performance
In the first set of simulations, we demonstrate that in the presence of $F$ malicious agents, indexed by $i\in\{0,\dots,F-1\}$, normal agents $i\in\{F,\dots,n-1\}$ in $\gamma$- and $(\gamma,\gamma)$-MERGs are guaranteed to achieve resilient consensus. Each normal agent $i$’s initial state $x_{i}[0]\in\mathbb{R}$ is randomly generated on the interval [1000,1000][-1000,1000]. At time step $t\geq 0$, a malicious agent $i$ transmits an adversarial value: $1080\cos(t/5)$ if $i$ is even, and $1080\sin(t/5)$ if $i$ is odd.
We consider networks with $n=49$ and $n=50$ agents in all simulations. Maximum $r$-robustness any network with $49$ and $50$ agents achieve is $r=\lceil\frac{n}{2}\rceil=25$-robustness. Thus, using Theorem 1, we construct $25$-MERGs with $49$ and $50$ nodes. Note these graphs achieve $25$-robustness with the fewest edges possible for a given number of nodes, and it tolerates at most $F=12$-local malicious agents. Thus, we run consensus simulations on these network graphs with $12$ malicious agents, and the results are shown in Figure 5 (a) and (b). Similarly, the maximum possible $(r,s)$-robustness any network with $49$ and $50$ agents achieves is $(25,25)$-robustness. Thus, we use Theorem 2 to construct $(25,25)$-MERGs with $49$ and $50$ agents. These graphs tolerate up to $F=24$-total malicious agents. We run consensus simulations on these network graphs with $24$ malicious agents, and the results are shown in Figure 5 (c) and (d). As shown in Figure 5, both $\gamma$- and $(\gamma,\gamma)$-MERGs allow all normal agents to successfully achieve consensus within the convex hull of their initial values through the W-MSR algorithm with $F$ malicious agents.
8.2 Consensus with an Edge Removal
These $\gamma$- and $(\gamma,\gamma)$ graphs achieve their respective maximum robustness levels using the fewest edges possible for a given number of nodes. In other words, removing any edge reduces their robustness levels at least by one. To validate this minimalistic property, we compare the consensus performance of the original $\gamma$- and $(\gamma,\gamma)$-MERGs with that of modified versions in which a single edge has been removed. For each graph, comparison is done using identical initial states for the normal agents and in the presence of $F$ misbehaving agents.
8.2.1 Simulations with $\gamma$-MERGs
We construct $5$-MERGs with $n=9$ and $n=10$ nodes as shown in Figure 2. Note these graphs achieve $5$-robustness with the fewest edges possible for a given number of nodes, and they tolerate at most $F=2$-local Byzantine agents. We set agents $i\in\{0,1\}$ as Byzantine, and agents $i\in\{2,\dots,n-1\}$ as normal. Each normal agent $i$ is initialized with a state $x_{i}[0]\in\mathbb{Z}_{\geq 0}$ sampled uniformly at random from the interval $[0,49]$ for $i\in\{2,3,4\}$ and from $[50,100]$ for $i\in\{5,\dots,n-1\}$. For all $t\geq 0$, all Byzantine agents adversarially share a value of $100$ with agents $i=0,\dots,\lceil\frac{n}{2}\rceil$ and of 100-100 with other remaining agents.
Figure 6 (a)-(b) and (e)-(f) illustrate the consensus results without and with edge removal. We have removed the edge $(3,8)$ for $n=9$ and $(4,9)$ for $n=10$ from the $5$-MERGs shown in Figure 2. In the $5$-MERGs, normal agents successfully reach resilient consensus (shown in Figure 6 (a) and (b)). However, removing the edge from $5$-MERG reduces its robustness to $r=4$, which is insufficient to guarantee resilient consensus in the presence of $F=2$ Byzantine agents. Thus, normal agents in the new graphs fail to reach resilient consensus (as shown by the red arrows in Figure 6 (e) and (f)).
8.2.2 Simulations with $(\gamma,\gamma)$-MERGs
We use Theorem 2 to construct $(5,5)$-MERGs with $n=9$ and $n=10$ agents, as shown in Figure 4. These graphs tolerate up to $F=4$-total malicious agents. We set agents $i\in\{0,\dots,3\}$ as malicious, and agents $i\in\{4,\dots,n-1\}$ as normal. Each normal agent $i$ is initialized with a state $x_{i}[0]\in\mathbb{R}$ sampled uniformly at random from the interval from $[0,49]$ for agents $i\in\{4,\dots,7\}$ and from $[50,100]$ for the other normal agents. Malicious agents share adversarial values: agent $3$ transmits 0 to its neighbors at all times, while the remaining malicious agents transmit $100$ to their neighbors at all times.
Figure 6 (c)-(d) and (g)-(h) illustrate the results without and with edge removal. We have removed the edge $(7,8)$ for $n=9$ and $(5,8)$ for $n=10$ from the $(5,5)$-MERGs visualized in Figure 4. In these $(5,5)$-MERGs, normal agents successfully reach resilient consensus (shown in Figure 6 (c) and (d)). However, removing the edge from each $(5,5)$-MERG reduces its robustness to at most $(5,4)$-robustness, which is insufficient to guarantee resilient consensus in the presence of $F=4$ malicious agents. As a result, normal agents fail to reach resilient consensus in the modified graphs (as shown by the red arrows in Figure 6 (g) and (h)).
9 Conclusion
This paper investigates and presents the necessary conditions and topological structures of $r$- and $(r,s)$-robust graphs. These results are derived purely from the definitions of robustness, providing a general understanding of what these robustness levels require. We establish the tight lower bounds on the edge counts for graphs with any number of nodes to achieve maximum robustness. Then, we develop further from these results by constructing two classes of graphs of any number of nodes that achieve maximum $r$- and $(r,s)$-robustness with minimal sets of edges. For future work, we aim to extend our results to arbitrary robustness levels.
10 Appendix
Here, the proof for Lemma 3 is provided below, accompanied by Figure 7 for illustrative support.
Figure 7: These figures demonstrate the first two iterations of the first part in the proof of Lemma 3. Throughout figures (a) to (d), we maintain $|\mathcal{S}_{1}|=\gamma-1$ and $|\mathcal{S}_{2}|=\gamma+1$. Figures (a) and (b) illustrate the first iteration. Here, $\mathcal{S}_{2}$ contains $\gamma+1$ nodes including $k$-clique $\mathcal{C}_{k,1}=\{v_{1},\cdots,v_{k}\}$ (colored in green), while $\mathcal{S}_{1}$ contains the remaining $\gamma-1$ nodes. Since $\mathcal{S}_{1}$ has to be $\gamma$-reachable, there exists a node $u_{1}\in\mathcal{S}_{1}$ that has edges with at least $\gamma$ nodes in $\mathcal{S}_{2}$. Figure (a) shows the case scenario where the yellow node $u_{1}$ does not have an edge with $v_{1}$ (highlighted in yellow). Then, $u_{1}$ forms a $k$-clique $\mathcal{C}_{k,2}$ with $\mathcal{C}_{k,1}\setminus\{v_{1}\}$. Consequently, in any future iteration, if any node $u_{i}\in\mathcal{S}_{1}$, $i\neq 1$, connects to at least $\gamma$ nodes in $\mathcal{S}_{2}$ but not to $v_{1}$, it will form a $(k+1)$-clique with $\mathcal{C}_{k,2}$. Then, $u_{1}$ gets swapped with a node in $S_{2}\setminus\mathcal{C}_{k,1}$, as shown in Figure (b). Figures (c) and (d) present the second iteration. Figure (c) presents the case scenario where $u_{2}$ only has $\gamma$ edges and does not have an edge with $v_{2}$ (highlighted in red). As before, if any node $u_{i}\in\mathcal{S}_{2}$, $i\neq 2$, connects to $\gamma$ nodes while missing an edge to $v_{2}$ in future iteration, it will form a $(k+1)$-clique. Then, $u_{2}$ gets swapped with a node in $\mathcal{S}_{2}\setminus U^{\prime}$ where $U^{\prime}=\mathcal{C}_{k,1}\cup\mathcal{C}_{k,2}$, as shown on Figure (d). The process continues, maintaining $\gamma$-reachability of $\mathcal{S}_{1}$, until a $\big(\lfloor\frac{\gamma+3}{2}\rfloor\big)$-clique is formed.
Proof.
Since $\mathcal{G}$ is $\gamma$-robust, it holds that for any pair of disjoint, non-empty subsets $\mathcal{S}_{1}$, $\mathcal{S}_{2}\subset\mathcal{V}$, at least one of them is $\gamma$-reachable. The proof is divided into two parts. In the first part, we will inductively show our argument by constructing two subsets $\mathcal{S}_{1},\mathcal{S}_{2}\subset\mathcal{V}$ such that $\mathcal{G}$ is $\gamma$-robust and contains a $c$-clique where $2\leq c\leq\lfloor\frac{\gamma+3}{2}\rfloor$. Then, in the second part of the proof, we will show that $\mathcal{G}$ must contain a $\big(\lfloor\frac{\gamma+4}{2}\rfloor\big)$-clique through some reasoning.
First Part: For the induction step, we fix $|\mathcal{S}_{1}|=\gamma-1$ and $|\mathcal{S}_{2}|=\gamma+1$, enforcing $\mathcal{S}_{1}$ to be $\gamma$-reachable. Hence we have the base case of a $2$-clique. Now, we will iteratively swap nodes between $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ to argue some of the edges are necessary for $\mathcal{G}$ to satisfy Definition 6.
WLOG, let $\mathcal{S}_{2}$ contain $\gamma+1$ nodes including $k$-clique $\mathcal{C}_{k,1}=\{v_{1},\cdots,v_{k}\}$ (green in Figure 7), where $k\geq 2$, and let $\mathcal{S}_{1}$ contain the remaining $\gamma-1$ nodes. We also denote $U_{1}=U_{1}^{\prime}=\mathcal{C}_{k,1}$. Since $\mathcal{S}_{1}$ is $\gamma$-reachable, $\exists u_{1}\in\mathcal{S}_{1}$ (yellow in Figure 7) such that $|\mathcal{N}_{u_{1}}\cap\mathcal{S}_{2}|=|\mathcal{N}_{u_{1}}\setminus\mathcal{S}_{1}|\geq\gamma$. There are two cases:
  1. 1.
    $u_{1}$ has edges with all $\gamma+1$ nodes in $\mathcal{S}_{2}$, or $\gamma$ nodes but not with a node in $\mathcal{S}_{1}\setminus U_{1}$.
  2. 2.
    $u_{1}$ has edges with $\gamma$ nodes but not with $v_{1}\in U_{1}$ (highlighted in yellow in Figure 7).
If 1), $u_{1}$ forms a $(k+1)$-clique with $\mathcal{C}_{k,1}$. If 2), $u_{1}$ forms another $k$-clique $\mathcal{C}_{k,2}$ with $U_{1}^{\prime}\setminus\{v_{1}\}$. Note that $u_{1}$ has edges with all nodes in $\mathcal{S}_{2}\setminus\{v_{1}\}$. Now, if 1), we have a $(k+1)$-clique and done. If 2), let $U_{2}=\mathcal{C}_{k,1}\cap\mathcal{C}_{k,2}=\{v_{2},\cdots,v_{k}\}$ and $U_{2}^{\prime}=\mathcal{C}_{k,1}\cup\mathcal{C}_{k,2}=\{v_{1},\cdots,v_{k},u_{1}\}$. We also swap $u_{1}$ with any of the nodes in $\mathcal{S}_{2}\setminus U_{1}^{\prime}$ to get $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ shown on Figure 7 (b).
Then, since $\mathcal{S}_{1}$ has to be $\gamma$-reachable, $\exists u_{2}\in\mathcal{S}_{1}$ (red in Figure 7) such that $|\mathcal{N}_{u_{2}}\cap\mathcal{S}_{2}|\geq\gamma$. Again, there are two cases:
  1. 1.
    $u_{2}$ has edges with all $\gamma+1$ nodes in $\mathcal{S}_{2}$, or $\gamma$ nodes but not with a node in $\mathcal{S}_{2}\setminus U_{2}$
  2. 2.
    $u_{2}$ has edges with $\gamma$ nodes but not with $v_{2}\in U_{2}$ (highlighted in red in Figure 7).
If 1), $u_{2}$ forms a $(k+1)$-clique with $\mathcal{C}_{k,j}$ for some $j\in\{1,2\}$. If 2), $u_{2}$ forms another $k$-clique $\mathcal{C}_{k,3}$ with $U_{2}^{\prime}\setminus\{v_{1},v_{2}\}=\{v_{3},\cdots,v_{k},u_{1}\}$. Also note that $u_{2}$ has edges with all nodes in $\mathcal{S}_{2}\setminus\{v_{2}\}$ including $u_{1}$. Now, if 1), we have a $(k+1)$-clique and done. If 2), we update $U_{3}=\bigcap\limits_{j=1}^{3}\mathcal{C}_{k,j}=\{v_{3},\cdots,v_{k}\}$ and $U_{3}^{\prime}=\bigcup\limits_{j=1}^{3}\mathcal{C}_{k,j}=\{v_{1},\cdots,v_{k},u_{1},u_{2}\}$. We then swap $u_{2}$ with any node in $\mathcal{S}_{2}\setminus U_{2}^{\prime}$, continuing the process.
Likewise, unless we get case 1), we keep encountering the second case illustrated in the previous steps. To generalize the procedure, let $u_{n}\in\mathcal{S}_{1}$ such that $|\mathcal{N}_{u_{n}}\cap\mathcal{S}_{2}|\geq\gamma$, $n\in\{1,\cdots,k\}$. Then, $U_{n}=\bigcap\limits_{j=1}^{n}\mathcal{C}_{k,j}=\{v_{n},\cdots,v_{k}\}$ and $U_{n}^{\prime}=\bigcup\limits_{j=1}^{n}\mathcal{C}_{k,j}=\{v_{1},\cdots,v_{k},u_{1},\cdots,u_{n-1}$}. There are two cases:
  1. 1.
    $u_{n}$ has edges with all $\gamma+1$ nodes in $\mathcal{S}_{2}$, or $\gamma$ nodes but not with a node in $\mathcal{S}_{2}\setminus U_{n}$.
  2. 2.
    $u_{n}$ has edges with $\gamma$ nodes in $\mathcal{S}_{2}$ but not with $v_{n}\in U_{n}$.
If 1), $u_{n}$ forms a $(k+1)$-clique with $\mathcal{C}_{k,j}$ for any $j\in\{1,\cdots,n\}$. If 2), $u_{n}\in\mathcal{S}_{1}$ forms another $k$-clique $\mathcal{C}_{k,n+1}$ with $U_{n}^{\prime}\setminus\{v_{1},\cdots,v_{n}\}$. Now, if 1), we have a $(k+1)$-clique and done. If 2) we swap $u_{n}$ with any node in $\mathcal{S}_{2}\setminus U_{n}^{\prime}$ to continue the procedure.
If we continue encountering the case 2) for $k$ times, since $\mathcal{S}_{1}$ is $\gamma$-reachable, $\exists u_{k+1}\in\mathcal{S}_{1}$ such that $|\mathcal{N}_{u_{k+1}}\cap\mathcal{S}_{2}|\geq\gamma$. Then, there are two cases:
  1. 1.
    $u_{k+1}$ has edges with all $\gamma+1$ nodes in $\mathcal{S}_{2}$.
  2. 2.
    $u_{k+1}$ has edges with $\gamma$ nodes in $\mathcal{S}_{2}$ but not with a node in $\mathcal{S}_{2}\setminus U_{k+1}$ where $U_{k+1}=\bigcap\limits_{j=1}^{k+1}\mathcal{C}_{k,j}=\emptyset$.
In both cases, $u_{k+1}$ forms a $(k+1)$-clique with $\mathcal{C}_{k,j}$ for any $j\in\{1,\cdots,k+1\}$, showing a $(k+1)$-clique must exist. Also note that $U_{k+1}^{\prime}=\{v_{1},\cdots,v_{k},u_{1},\cdots,u_{k}\}\subseteq S_{2}$. Since (i) at worst $u_{1},\cdots,u_{k}$ as well as $v_{1},\cdots,v_{k}$ need to be in $\mathcal{S}_{2}$ for a $(k+1)$-clique to be formed and (ii) $|\mathcal{S}_{2}|=\gamma+1$, $k\leq\lfloor\frac{\gamma+1}{2}\rfloor$. That means $\mathcal{G}$ contains a $\big(\lfloor\frac{\gamma+3}{2}\rfloor\big)$-clique, which completes the first part of the proof.
Second Part: Continuing from the previous paragraph, at $k=\lfloor\frac{\gamma+1}{2}\rfloor$ and for even values of $\gamma$, at worst the case 2) mentioned above is repeated $\frac{\gamma}{2}$ times. Then, $\mathcal{S}_{2}$ contains $\mathcal{K}=\{u_{1},\cdots,u_{\frac{\gamma}{2}}\}$, $\mathcal{C}_{k,1}=\{v_{1},\cdots,v_{\frac{\gamma}{2}}$}, and a node $p$ not in $\mathcal{K}$ and $\mathcal{C}_{k,1}$. Note both $\mathcal{K}$ and $\mathcal{C}_{k,1}$ disjointly form a $\big(\frac{\gamma}{2}\big)$-clique. We know that a node $u_{m}\in\mathcal{S}_{1}$, $m=\frac{\gamma+2}{2}$, must form a $m$-clique $\mathcal{Q}$ with either $\mathcal{K}$ or $\mathcal{C}_{k,1}$. We swap $u_{m}\in\mathcal{S}_{1}$ with $p\in\mathcal{S}_{2}$. Now let $\mathcal{P}=\mathcal{S}_{2}\setminus\mathcal{Q}$. Note that nodes in $\mathcal{P}$ form a $(\frac{\gamma}{2})$-clique such that $\mathcal{P}\cap\mathcal{Q}=\emptyset$. Since $\mathcal{S}_{1}$ is $\gamma$-reachable even after $u_{m}$ is swapped out, $\exists u_{m+1}\in\mathcal{S}_{1}$ such that $|\mathcal{N}_{u_{m+1}}\cap\mathcal{S}_{2}|\geq\gamma$. There are two cases:
  1. 1.
    $u_{m+1}$ has edges with all $\gamma+1$ nodes in $\mathcal{S}_{2}$, or $\gamma$ nodes but not with a node in $\mathcal{P}$.
  2. 2.
    $u_{m+1}$ has edges with $\gamma$ nodes in $\mathcal{S}_{2}$ but not with a node in $\mathcal{Q}$.
If 1), $\mathcal{Q}\cup\{u_{m+1}\}$ forms a $\big(\frac{\gamma+4}{2}\big)$-clique. If 2), $\mathcal{P}\cup\{u_{m+1}\}$ forms a $\big(\frac{\gamma+2}{2}\big)$-clique $\mathcal{P}^{\prime}$ such that $\mathcal{P}^{\prime}\cap\mathcal{Q}=\emptyset$. In this case, WLOG, let $\mathcal{P}^{\prime}\subset\mathcal{S}_{1}$ and $\mathcal{Q}\subset\mathcal{S}_{2}$ such that $|\mathcal{S}_{1}|=|\mathcal{S}_{2}|=\gamma$. Since either $\mathcal{S}_{1}$ or $\mathcal{S}_{2}$ is $\gamma$-reachable, all of the nodes in either $\mathcal{P}^{\prime}$ or $\mathcal{Q}$ have edges with one additional node in $\mathcal{S}_{2}$ or $\mathcal{S}_{1}$ respectively, forming a $\big(\frac{\gamma+4}{2}\big)$-clique. Thus, $\mathcal{G}$ must contain a $\big(\frac{\gamma+4}{2}\big)$-clique for even $\gamma$. Since $\mathcal{G}$ must have a clique size of $\frac{\gamma+3}{2}$ and $\frac{\gamma+4}{2}$ for odd and even $\gamma$ respectively, $\mathcal{G}$ must have a $\big(\lfloor{\frac{\gamma+4}{2}}\rfloor\big)$-clique. ∎